### v2

# The Compilation and Validation of the Spectroscopic Redshift Catalogs for the DESI-COSMOS and DESI-XMMLSS Fields

Abstract
Over several dedicated programs that include targets beyond the main cosmological samples, the Dark Energy Spectroscopic Instrument (DESI) collected spectra for 311,102 unique objects in two fields centered on the COSMOS and XMM-LSS fields. In this work, we develop spectroscopic redshift robustness criteria for those spectra, validate these criteria using visual inspection, and provide two custom Value-Added Catalogs with our redshift characterizations. With these criteria, we reliably classify 217,037 galaxies below $z<1.6$, 9,688 quasars and 35,505 stars. As a critical element in characterizing the selection function, we provide the description of 70 different algorithms that were used to select these targets from imaging data. To facilitate joint imaging/spectroscopic analyses, we provide row-matched photometry from the Dark Energy Camera, Hyper-Suprime Cam, and public COSMOS2020 photometric catalogs. Finally, we demonstrate example applications of these large catalogs to photometric redshift estimation, cluster finding, and completeness studies.
1 Introduction
The Dark Energy Spectroscopic Instrument (DESI) is a Stage-IV dark energy spectroscopic survey currently being used to explore the cosmological model through measurements of large scale structure (DESI Collaboration et al., 2016, 2025a, 2025b). The instrument is designed to measure the baryonic acoustic oscillation (BAO) feature at sub-percent precision over each of three redshift intervals from $0<z<2.5$. Making the measurement at such high precision requires measuring redshifts of tens of millions of galaxies and quasars (DESI Collaboration et al., 2024). With a high fiber count and wide field of view to meet this requirement, DESI is also an ideal instrument for building customized spectroscopic samples at very high surface density over a few square degrees.
The COSMOS and XMM-LSS fields are two equatorial regions, with a legacy of multiband-photometry and spectroscopy. The COSMOS Survey (Scoville et al., 2007), covering 2 deg2, pioneered deep photometry over many bandpasses to facilitate a broad range of astronomical studies. The COSMOS2020 data release (Weaver et al., 2022) includes imaging programs that span radio to X-ray wavelengths from the following facilities: Very Large Array (Schinnerer et al., 2010), Max-Planck Millimeter Bolometer Array (Bertoldi et al., 2007; Scott et al., 2008), Spitzer (Sanders et al., 2007), Visible and Infrared Survey Telescope for Astronomy (McCracken et al., 2012), unWISE (Schlafly et al., 2019), Subaru, Hubble Space Telescope, Mayall, Blanco, CFHT, (Taniguchi et al., 2015; Koekemoer et al., 2007; Capak et al., 2007), GALEX (Zamojski et al., 2007), XMM-Newton (Cappelluti et al., 2009) and Chandra (Civano et al., 2016). The public COSMOS catalogs also includes a large collection of photometric redshift, stellar mass, and star formation rate estimates. The latest COSMOS2025 data release (Shuntov et al., 2025) includes imaging from JWST NIRcam over the central 0.54 deg2 of the COSMOS field. The XMM-LSS field also benefits from an overlap in multi-band photometry and spectroscopic data. The multi-band photometry spans from radio to X-ray from the following facilities: GMRT (Smolčić et al., 2018), LoFAR (Hale et al., 2019), HerMES (Oliver et al., 2012), SWIRE (Lonsdale et al., 2003), Dark Energy Survey(Abbott et al., 2018), Hyper Suprime-Cam (Aihara et al., 2022), Galex (Zamojski et al., 2007), XMM-SERVS and XXL (Chen et al., 2018; Pierre et al., 2016).
The COSMOS and XMM-LSS deep, multiband-photometry has been complemented by various spectroscopic programs including SDSS, BOSS, and eBOSS (York et al., 2000; Eisenstein et al., 2011; Dawson et al., 2013, 2016; Blanton et al., 2017), VLT/VIMOS (Lilly et al., 2007), OzDES (Childress et al., 2017), DEVILS (Davies et al., 2018), MMT/Binospec (Fabricant et al., 2019), C3R2 DR3 (Stanford et al., 2021), VIPERS (Scodeggio et al., 2018), DEIMOS (Hasinger et al., 2018), LEGA-C DR2 (Straatman et al., 2019), MOSDEF (Kriek et al., 2015), VVDS (Le Fèvre et al., 2015), and DEEP2 (Newman et al., 2013). Recently, the COSMOS spectroscopic redshift compilation [^1] https://github.com/cosmosastro/speczcompilation was publicly released to summarize all published redshifts within the 2 deg2 COSMOS field (Khostovan et al., 2025).
The wealth of existing photometric and spectroscopic data has supported a large set of astrophysical studies such as the calibration of photometric redshifts (e.g., Myles et al., 2021) for cosmic shear and weak lensing measurements (e.g., Abbott et al., 2022). In addition, these data sets have allowed estimates of the stellar mass function (e.g., Davidzon et al., 2017), studies of the integrated stellar, gas, dust, and dark matter properties for high redshift galaxies (e.g., Sillassen et al., 2024), and constraints on the UV-luminosity function and cosmic star formation rate density (e.g., Donnan et al., 2023). Other examples include the confirmation of proto-clusters and cluster members (e.g., McConachie et al., 2022) and spectroscopic confirmation of Ly-$\alpha$ emitting galaxies at $z>7$ to probe the high redshift universe (e.g., Cooper et al., 2024).
While published data have supported a range of science, the existing spectroscopic samples come from multiple instruments with varying wavelength coverage and robustness in redshift classification. In addition, the algorithms for selecting targets from imaging data are rarely defined, leaving uncertainty in the selection function. DESI, with well-described target selection algorithms and consistent data processing methodology, provides an opportunity for a well-calibrated spectroscopic sample to complement the existing photometry.
DESI dedicated 50 hours to spectroscopy in a 8 deg2 region centered on the COSMOS field and 17 hours in a 8 deg2 region centered on the XMM-LSS field. With its 5000-fiber multiplex, when also accounting for its main survey, DESI produced 233,247 and 74,990 unique spectra over 16 deg2 centered on the COSMOS and XMM-LSS fields, respectively. We define the DESI-COSMOS ($\alpha=150.1^{\circ}$, $\delta=2.2^{\circ}$) and DESI-XMMLSS ($\alpha=35.6^{\circ}$, $\delta=-4.8^{\circ}$)fields in Figure 1 to cover 16 deg2, ensuring complete overlap with existing and future deep imaging areas. Because the DESI-COSMOS field has more dedicated observation time, that field has a higher density sample over the central 8 deg2 region.
Figure 1: The spectroscopic coverage of the DESI Value-Added catalogs. The boundaries of the plot enclose the 16 deg2 DESI-COSMOS and DESI-XMMLSS fields. The colorbar depicts the number of objects per arcmin2. Left: Over-plotted on the DESI-COSMOS field are overlapping deep drilling fields for COSMOS2020, COSMOS2025, LSST and HSC COSMOS-Deep/Ultradeep fields. Right: Over-plotted on the DESI-XMMLSS field is the LSST deep drilling field and combined HSC Deep XMM-LSS and Ultradeep SXDS fields.
In these two fields, DESI’s five primary samples were supplemented with a variety of customized target samples. The five primary targeting programs include Luminous Red Galaxies (LRG; Zhou et al., 2023), the Bright Galaxy Survey (BGS; Hahn et al., 2023), the Milky Way Survey (MWS; Cooper et al., 2023), Emission Line Galaxies (ELG; Raichoor et al., 2023) and Quasars (Chaussidon et al., 2023). The supplementary programs included samples at a low surface density, allocated spare fibers during the primary program, and samples at a higher surface density that required dedicated observations with customized fiber assignments.
In this work, we present a custom evaluation of all spectra within the COSMOS and XMM-LSS fields. We modify the redshift quality cuts beyond those that were informed by visual inspection for DESI galaxies (Lan et al., 2023) and quasars (Alexander et al., 2023). The modified quality cuts increase the redshift purity for the main samples while minimally reducing the completeness in these fields. We use these new criteria to develop an algorithm for determining reliable redshifts for the wealth of supplemental programs within these two fields. In total, over 32 deg2, we find robust spectroscopic classification of 217,037 galaxies below $z<1.6$, 9,688 quasars and 35,505 stars. In addition, several dedicated programs acquired spectra for potential Lyman-break and Lyman-alpha emission galaxy targets (Ruhlmann-Kleider et al., 2024; White et al., 2024) with full redshift performance to be reported in the future. We exclude these targets from our study, as the automated redshift classification is still in development.
Complementing this study, we are publicly releasing a Value-Added Catalog that provides redshift and quality estimates for galaxies and quasars and radial velocities and derived parameters for stars. These catalogs also include photometric measurements from DECaLs DR9222https://www.legacysurvey.org/dr9/description/, Hyper Suprime-Cam PDR3 Wide/Ultra Deep (Aihara et al., 2022), DECam DR10333https://www.legacysurvey.org/dr10/description/, COSMOS2020 (Weaver et al., 2022) and the Merian medium-band survey from the Victor M. Blanco telescope (Luo et al., 2024).
In addition to the legacy value, this catalog has potential to support future work such as photometric observations from the Vera C. Rubin Observatory. With its 9.6 deg2 field of view, the Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST) will produce a large photometric sample spanning 20,000 deg2 of the sky (LSST Science Collaboration et al., 2009). This survey will include frequent, dedicated observations in the COSMOS and XMM-LSS fields (Gris et al., 2024), which are well-matched to the DESI 3-degree-diameter field of view. Of the many science goals within LSST, weak lensing and cosmic shear serve as clear examples of science drivers that will benefit from having spectroscopic observations in these two dedicated fields.
Here, we present the samples, classification scheme, and redshift robustness statistics for the Value-Added Catalog. In Section 2, we describe the DESI target selection algorithms. In Section 3, we describe the observations of the COSMOS and XMM-LSS fields, the automated redshift classification pipeline, redshift consistency tests, spectral modeling, and visual inspection procedures. In Section 4 we present the quality cuts for each of the main target classes and how those cuts inform the quality assessment for each of the supplemental programs. We also present the results of consistency and robustness checks. In Section 5, we describe the Value-Added Catalog of the DESI spectroscopic redshifts in the COSMOS and XMM-LSS fields. Finally, in Section 6 we provide examples on the utility of these spectroscopic redshifts for other cosmological studies and provide a summary in Section 7. The Value-Added Catalogs contain large heterogeneous samples of extragalactic and stellar sources. Due to the complexity of all the samples, we provide selection functions and redshift robustness criteria for each target class to enable various scientific studies. All magnitudes in the paper are in AB magnitude system (Oke & Gunn, 1983) and are corrected for Galactic extinction. The photometry included in the matched catalogs remains in the standard format, without corrections from Galactic extinction.
2 DESI Target Classes
The majority of targets chosen for DESI observation were selected from the DESI Legacy Imaging Surveys (Dey et al., 2019; Zou et al., 2017). Additional targets were chosen from Hyper Suprime Cam (HSC) PDR3 Wide/Ultra Deep, the Dark Energy Camera Deep observations, and Merian photometry. The selection procedure for targets is given in Myers et al. (2023).
DESI observations were planned in four different phases. The first phase of the DESI program was Target Validation (December 2020 to March 2021), hereafter referred to as SV1. SV1 consisted of deep observations of extended target selections for each of LRG, ELG, quasar, BGS and MWS targets (DESI Collaboration et al., 2024). SV1 data were then used to finalize the MAIN Survey target selections. The next major phase of the Survey Validation program was the One-Percent Survey, hereafter called SV3 (April – May 2021). Targets in this survey were selected by algorithms that closely represented those in the MAIN Survey that were assigned fibers for spectroscopic observation at a very high completeness. Planning for the MAIN Survey, included fiber assignments over 14,000 deg2 and is ongoing with the primary goal of constraining the nature of dark energy (DESI Collaboration et al., 2025c, d). Finally, customized observations were planned as special programs in dedicated fields.
The dark time targets (LRGs, ELGs, quasars) for the main survey were assigned fibers over a tiling pattern that yielded an average of roughly five observation opportunities per coordinate. The bright time targets (BGS and MWS) yielded an average of roughly three observation opportunities per coordinate. For SV1, SV3 and in the special programs, customized tile patterns provided more observation opportunities per coordinate and thus a higher surface density of targets that were assigned fibers In total 146 and 44 special dedicated tiles were assigned to COSMOS and XMM-LSS, respectively.
2.1 DESI Target Bitmasks
DESI targets are identified for spectroscopy via the algorithm outlined in Myers et al. (2023), where each object is assigned a bitwise number indicating the target class for which that object was targeted. DESINAME was introduced in 2024 to help identify objects that undergo multiple observations across DESI phases. DESINAME is particularly important as it serves as the main identifier for all targets across the catalog.
For each phase, the selection of targets was recorded via bitmasks. For each of SV1, SV3, and MAIN, there is a primary, secondary, MWS, and BGS bitmask, totaling 12 bitmasks. For targets in special programs, target information is recorded in text files external to the DESI target bitmasks as indicated by the DARK_TOO_HIP bit in the secondary mask.
Each target has an associated DESI_TARGET, DESI_SCND_TARGET, BGS_TARGET, and/or MWS_TARGET bit that indicates the targeting classes and subclass. In this scheme, the DESI_TARGET bitmask relays the top-level information including the class of primary targets (LRG, ELG, quasar, BGS and MWS) and where subclass information can be found in additional bitmasks. Subclass information for MWS targets can be found in MWS_TARGET and BGS subclass information can be found BGS_TARGET. Secondary subclass information can be found in DESI_SCND_TARGET. Prefixes for SV1 and SV3 are appended to the bitmask if an object was observed prior to the MAIN survey.
This bitmask structure leads to a selection algorithm that is easily recorded for each of the targets. The overall selection function is therefore well understood in both the COSMOS and XMM-LSS fields. In what follows, we describe the target selection philosophy for the primary, secondary and special program targets. We provide a full description of each target selection in Appendix A.
2.2 Primary Target Selection
The primary target classes of LRG, ELG, quasar, BGS and MWS targets comprise the majority of spectra produced by DESI. The philosophy for the primary DESI target selection is outlined below.
2.2.1 Luminous Red Galaxies (LRG)
The LRG sample is selected from imaging data based on a strong continuum break at 4000 Å. The sample is selected using $g,r,z$ optical bands with infrared photometry from the WISE $W1$ band with an emphasis on the rW1r-W1 color. This rW1r-W1 color serves as good proxy for redshift from $0.4<z<1.1$ for a passively evolving population with a strong 4000 Å  break (Zhou et al., 2023). This selection is intended to cover an approximately constant number density over as wide of a redshift range as possible.
The DESI LRG sample has a significantly higher target density and redshift range when compared to previous surveys such as eBOSS and SDSS (Zhou et al., 2023). In addition to the main selection of LRG targets, special programs were designed to increase the number density by almost an order of magnitude in the COSMOS field.
For all these samples, these objects are dominated by older stellar populations, yielding redder spectra with relatively strong metal absorption lines allowing us to measure robust redshifts.
2.2.2 Emission Line Galaxies (ELG)
The ELG sample is selected based on relatively blue colors in the grg-r vs. rzr-z plane. In the approximate redshift range from $0.6<z<1.6$, these colors signify the presence of a young stellar population. By focusing on this higher redshift range, this selection identifies a population of galaxies in a redshift regime where LRGs are too faint to be efficiently classified through DESI spectroscopy.
The DESI ELG sample has a target density about 10 times more dense than the ELG sample of eBOSS and stretches one magnitude fainter (Raichoor et al., 2023). Several special programs were designed to explore emission line galaxies in the COSMOS field, further increasing the number density.
These galaxies become very faint in continuum flux at higher redshift but commonly feature strong [OII] emission lines. The [OII] doublet $\lambda\lambda 3726,3279$ Å resulting from star formation (Madau & Dickinson, 2014) can be used to uniquely determine the spectroscopic redshift of the target.
2.2.3 Quasars
Quasars are extremely luminous active galactic nuclei powered by accretion onto supermassive black holes. Their intrinsic brightness makes them efficient tracers of large scale structure at redshifts $z>0.9$. Quasars are selected through a combination of optical (g, r, z) and IR colors using the $W1$ and $W2$ measurements from WISE. Due to the dusty torus, quasars are brighter in the mid-infrared at all redshifts compared to stars of similar optical magnitude and color. The DESI selection uses this unique feature as a powerful tool to reject stars from the quasar selection (Chaussidon et al., 2023).
The bulk of the quasar sample comes from the main target selection, with small secondary programs supplementing the complete quasar sample in the COSMOS and XMM-LSS fields.
Quasar spectra are characterized by multiple broad emission lines that extend from the UV to visible wavelengths in the rest frame. These broad emission lines provide a useful tool for robust spectroscopic classification.
2.2.4 Bright Galaxy Survey (BGS)
The BGS sample will be a collection of over 10 million galaxies spanning $0<z<0.4$. The goal of the BGS sample is to probe the epoch where dark energy is dominant (Hahn et al., 2023).
[…]
Similar to the LRG sample, the relatively low redshift and intrinsic brightness of these targets lead to a strong continuum that enables the detection of a plethora of absorption lines, prime for redshift classification.
2.2.5 Milky Way Survey (MWS)
Making use of DESI bright time, the primary Milky Way Survey program will include approximately ten million stars to a limiting magnitude of $r<19$ with the goal of providing radial velocities, metallicities, and other stellar parameters (Cooper et al., 2023). The main target selection algorithm covers three target categories, MAIN-BLUE, MAIN-RED and MAIN-BROAD, which in combination cover the full color–magnitude space within $16<r<19$ (Cooper et al., 2023). In addition, there are 23 stellar target classes spanning the observational programs in DESI and stars often appear as interlopers in other selections as well.
2.3 Secondary Targets
Secondary programs, designed by members of the DESI collaboration, extend beyond DESI’s primary cosmology science drivers. Across both the COSMOS and XMM-LSS fields, there are targets that span 32 secondary programs. Examples of these programs include point sources with extreme colors, low redshift dwarf galaxies, high proper-motion stars, supernova host galaxies, and a variety of other extragalactic and stellar sources. Examples of results from these programs include calibration of the Fundamental Plane relation (Said et al., 2024), relationships between radio emission and dust extinction in quasars (Fawcett et al., 2023), and much more. We provide a full description of the secondary target selections in Appendix A.
2.4 Special Programs
DESI conducted dedicated observations of high-value, high-density targets to explore options for future surveys and enrich the science from the COSMOS and XMM-LSS fields. For each of these programs, a set of dedicated tiles, each with a customized target list, was designed for observation. The observations of these tiles were prioritized to be completed in a limited time frame.
The extended selections for LRGs and ELGs mentioned above were conducted in these special programs. Other samples from these programs include samples for photometric redshift calibration of LSST weak lensing sources, LAEs and LBGs. Publications from these programs include LAE clustering studies (White et al., 2024), LBG target selection techniques (Ruhlmann-Kleider et al., 2024) and extracting the Lyman-$\alpha$ forest from LBG spectra (Herrera-Alcantar et al., in prep). There were six special programs in the COSMOS field and five in the XMM-LSS field completed in the first three years of the DESI survey. We provide a full description of the special target selections in Appendix A.
3 Spectroscopy
DESI is a multi-object spectroscopic instrument installed on the Mayall 4-m telescope at Kitt Peak National Observatory in Arizona. The instrument makes use of an 8 deg2 field of view and robotic fiber positioners to simultaneously capture spectra across a focal plane of 5000 fibers (DESI Collaboration et al., 2022a; Poppett et al., 2024; Miller et al., 2024). Ten petals of 500 fibers each connect to one of ten high-efficiency spectrographs with a spectroscopic wavelength coverage from 3600-9800 Å. The high throughput, efficient fiber positioning, and high fiber count allow DESI to capture thousands of spectra over a short period of time.
A standard observation in DESI consists of a 1000s effective exposure time (Schlafly et al., 2023). In addition to standard DESI observations, several hours of effective spectroscopic exposure time were dedicated to special tiles in the COSMOS field and the XMM-LSS field (DESI Collaboration et al., 2025e). To facilitate studies in these two regions and in neighboring areas, we produce two catalogs covering 16 deg2 each. The footprint of this study includes coverage from the standard tiles in addition to special tiles centered on the COSMOS and XMM-LSS fields. The fields covered in these catalogs hereafter are referred to as the DESI-COSMOS and DESI-XMMLSS fields.
These observations are processed through a data reduction pipeline which converts the raw CCD images to one-dimensional spectra (Guy et al., 2023) followed by redshift and spectroscopic classification. Using these redshift estimates, we then apply a spectral fitting algorithm that simultaneously models the spectrum’s continuum and emission lines. To further quantify automated redshift efficiency, we compare the pipeline redshift estimates across multiple observations of the object in different phases of DESI, when possible. We visually inspect all discrepancies in redshift to estimate the catastrophic failure rate and inform quality cuts customized to each target class.
3.1 The DESI Data Reduction Pipeline and Redshift Estimation
We summarize the data reduction pipeline which is described in full detail in Guy et al. (2023). The raw images are corrected for dark current, flat fields and bad pixels. Each science exposure undergoes wavelength calibration, and the spectroscopic Point Spread Function (PSF) is measured and used in a spectral extraction. Sky spectrum models are subtracted from spectra using fibers placed on empty areas of the sky (called “sky fibers”) and flux calibration is performed using appropriate stars. For each survey (SV1, SV3, MAIN, SPECIAL), a unique coadded spectrum is created for each object. That is, multiple coadds exist for the same object if it was observed in different surveys, allowing us to test the consistency of the redshift pipeline. With these co-added spectra, spectral classification and redshift estimates are performed.
DESI utilizes a Principal Component Analysis (PCA) fitting algorithm, called Redrock [^4] https://github.com/desihub/Redrock, to determine spectral class and redshift estimates for all spectra (Bailey et al., in prep; Anand et al., 2024). The algorithm uses a set of templates that represent broad spectral classes of galaxies, stars and quasars. The combination of these templates provides composite solutions constructed to best model each spectrum. Redrock fits these composite solutions across a complete range of redshift-parameter space, choosing the best fitting solution based on the lowest $\chi^{2}$ value following the philosophy in Bolton et al. (2012).
Redrock produces many fit parameters useful for assessing the quality of the template fit and redshift estimation. The $\chi^{2}$ can be used to determine the quality of the best template fit. Another fit statistic, $\Delta\chi^{2}$ is the difference between the $\chi^{2}$ of two neighboring fits (for Redrock, the best and next best fit). A high $\Delta\chi^{2}$ indicates that the best fit is highly preferred compared to the next best fit and indicates a higher confidence in the redshift estimate. A small $\Delta\chi^{2}$ indicates that there is little distinction between the first and second-best fit, leading to a lower confidence in the redshift estimate. All fits are given a ZWARN flag, indicating any issues in the observation or data reduction pipeline for that spectrum. Redrock also classifies the spectrum, attributing the spectral class, or SPECTYPE, to the template of the best fit.
For quasars specifically, we utilize two afterburner pipelines, MgII and QuasarNET (Green et al., 2025), to ensure proper spectral classification of these targets. The MgII afterburner tests for a broad MgII $\lambda 2800$ Å emission line, characteristic of qusar spectra. The method consists of fitting a Gaussian in a 250 Å window centered at the position of MgII line given by Redrock. We consider the MgII as a broad line if the improvement of $\chi^{2}$ is better than 16, the width of the Gaussian greater than 10 Å and the significance of the amplitude of the Gaussian greater than 3. The algorithm possibly changes the source classification but never modifies the redshift given by Redrock (Chaussidon et al., 2023).
[…]
3.2 FastSpecFit
Redshift estimation for objects with a low signal-to-noise ratio per pixel rely on prominent emission lines to accurately estimate a spectrum’s redshift. The strength of these emission lines can be an indicator of the success of the redshift estimation and used as a metric for redshift quality. To get a more accurate measurement of the strength of emission lines in these spectra, we use FastSpecFit [^5] https://github.com/desihub/fastspecfit(Moustakas et al., 2023), a physically-motivated algorithm that measures emission line strength and physical properties of a galaxy from its spectrum and photometry. FastSpecFit produces measurements for fluxes and uncertainties of emission lines such as, the [OII] $\lambda\lambda 3726,3729$ Å and [OIII] $\lambda\lambda 4959,5007$ Å lines. These measurements from FastSpecFit will be used to examine redshift quality for target classes with strong emission line properties.
3.3 Assessing Redshift Estimations
To assess the accuracy of the data reduction and Redrock pipeline, we compare objects across multiple observations. If the estimated redshifts for the same object are consistent, we assume that the redshift estimation is correct. In cases where we have discrepant redshift estimations, we assume one of the redshift estimations is incorrect.
We match objects that have the same DESINAME across the all programs, MAIN, SV3, and SV1. Before matching, we require ZWARN = 0 or 4 with a $\Delta\chi^{2}\geq 40$ threshold, informed by prior survey validation of the BGS sample (Lan et al., 2023), to remove lower quality spectra. ZWARN = 0 indicates there was no issue with the redshift fit, and ZWARN = 4 indicates that the $\chi^{2}$ of best fit is too close to that of second best ($\Delta\chi^{2}<9$. A ZWARN = 4 is non-problematic as we apply our own $\Delta\chi^{2}$ cuts, customized to each sample. We then compare the redshifts via Eq. 1 and Eq. 2.
Following the science requirements for DESI, a threshold for catastrophic failures is set at 1000 km/s (DESI Collaboration et al., 2022b). If the difference between reported redshifts of the same target in different programs is within 1000 km/s for galaxies and stars, the target is considered to have a robust and consistent redshift:
$\Delta v=\frac{|z_{i}-z_{j}|*c}{(1+z_{i})}\leq 1000~\rm{km\,s^{-1}}$ (1)
If the difference between reported redshifts is within 3000 km/s for quasars, the target is considered to have a robust and consistent redshift:
$\Delta v=\frac{|z_{i}-z_{j}|*c}{(1+z_{i})}\leq 3000~\rm{km\,s^{-1}}$ (2)
We find a high consistency in redshift estimation of better than $\sim$99.3% across programs for both the DESI-COSMOS and DESI-XMMLSS field as seen in Table 1. Examining the discrepant pairs allows us to identify failure modes and better understand catastrophic failures.
Survey # of Matches Catastrophic Failure Rate
COSMOS
MAIN-SV3 70250 0.35%
MAIN-SV1 14094 0.72%
SV3-SV1 20719 0.41%
XMM-LSS
MAIN-SV1 10871 0.37%
Table 1: Cross matched objects between DESI programs. The catastrophic failure rate is the implied rate of failure for all objects with ZWARN = 0 or 4 and $\Delta\chi^{2}\geq 40$.
3.4 Visual Inspection
In cases where measurements disagree, we visually inspect the individual spectra to investigate which are catastrophic failures. The goals of visual inspection of are to: (1) assess the DESI redshift estimation pipeline; and (2) investigate catastrophic failure modes of spectra to inform quality cuts. The visual inspection (VI) procedure is conducted through the Prospect [^6] https://github.com/desihub/prospect spectral visualization tool.
The standardized process for inspecting DESI spectra is outlined in Lan et al. (2023) and is summarized here:
  1. 1.
    Each spectrum has 2 visual inspectors.
  2. 2.
    Each inspector reports four key features:
    • Redshift
    • A quality flag where spectra are assigned 1 of 5 discrete values ranging from 0-4. 4 denotes the highest quality a spectrum can receive and 0 denotes the lowest score.
    • Type of source: galaxy, star or quasar
    • Any anomalies, systematic features, or artifacts within the spectrum
    • Additional comments regarding the spectra (e.g., two objects in the spectrum)
  3. 3.
    The outputs of the two inspectors are analyzed and compared. For each spectrum, the VI report must satisfy the following conditions:
    • the SPECTYPE from both inspectors is the same;
[…]
    • the difference between VI redshift is within 1000 km s-1.
  4. 4.
    If any of these conditions are not met, a third inspector inspects the spectrum and determines the final VI results. This third inspector might be one of the original inspectors.
The criteria for the quality of each spectra, denoted by a discrete value 0-4 is as follows (Lan et al., 2023):
  • Quality 4: confident classification based on clear two or more spectral features (e.g., spectra with multiple absorption lines or emission lines);
  • Quality 3: probable classification with at least one secure spectral feature and continuum or many weak spectral features;
  • Quality 2: possible classification with one strong spectral feature;
  • Quality 1: unlikely classification with some signal but no distinct feature;
  • Quality 0: no meaningful signal;
With the above procedure and criteria, a final VI redshift with an average VI quality $\geq$ 2.5 in considered a secure VI redshift.
In addition to the visual inspection on objects with discrepant redshifts, we spot check objects that have multiple redshift estimations across DESI without any quality cuts. Due to the size of this sample, systematically visually inspecting over 3,000 spectra is not tractable and spot checks were simply used to inform quality cuts.
4 Redshift Quality Assessment
Using the trends identified through visual inspection of objects with failed redshift estimations, we develop quality cuts to ensure robust redshift estimates for each target class. For all objects, we only include spectra with $\texttt{ZWARN}=0$ or $\texttt{ZWARN}=4$.
We find for the BGS and LRG samples, failures in redshift estimation stem from spectra with low SNR per pixel. ELG spectra without strong [OII] emission lead to poor redshift estimates as well. Line confusion for broad-line emission in quasars leads to incorrect classification and redshift failures. The primary target classes then serve as models for developing customized quality cuts for all of the secondary and special program targets. We use the LRG sample as the archetype for strong-continuum, high SNR per pixel spectra and the ELG sample for strong emission line spectra. All targets spectroscopically confirmed as quasars or targeted as a quasar are run through the quasar afterburners. For all spectra classified as stars, we simply report derived stellar parameter measurements from the RVSpecfit pipeline (Koposov et al., 2025, 2011).
4.1 Bright Galaxy Survey
The BGS selection was developed to prioritize objects with a strong continuum and therefore a high SNR across the spectrum. We find objects with a weak continuum have poor redshift estimates while objects with a high SNR continuum yield accurate redshift estimates.
Before any quality cuts are assigned, we assign all BGS spectra with a spectroscopic classification of ‘STAR’ or ‘QSO’ to be evaluated with the stellar and quasar quality assessments, respectively. We also impose a hard redshift cut of $z<1.5$ as there are known failure modes for classifications past this redshift.
Figure 2: Statistics of the BGS sample in the COSMOS field. Left: Distribution of the BGS sample in $\chi^{2}$ as a function of $\Delta\chi^{2}$. A large fraction of objects cluster in the locus around high $\Delta\chi^{2}$ and $\chi^{2}\sim 8000$. The red data points indicate visually inspected BGS spectra with incorrect redshifts. The red dashed line displays the $\chi^{2}$ and $\Delta\chi^{2}$ quality cuts. Spectra in the upper right are assumed to have robust redshifts. Right: The number of BGS spectra with robust redshifts in bins of $\Delta z=0.05$. The total BGS sample (black) is split into a BGS Bright sample ($r<19.5$, blue) and a BGS Faint sample (orange), which extends the selection to fainter galaxies.
[…]
Turning our attention to catastrophic redshift failures identified from visual inspection, we find poor redshift estimates at low $\Delta\chi^{2}$. This motivates a requirement that spectra have $\Delta\chi^{2}>40$ in order to be considered a robust redshift.
Spectra that have $\chi^{2}<7000$ tend to have a large fraction of their pixels flagged in the data reduction. We find that all of the pixels from one or more cameras were flagged, indicating at least one camera malfunctioned or was corrupted during observation. We therefore require spectra to have a $\chi^{2}>7000$ in order to be considered a robust redshift.
Spectra with high $\chi^{2}$ yield a significant fraction of catastrophic failures in the regime where $\Delta\chi^{2}<100$. A combination of low $\Delta\chi^{2}$ and high $\chi^{2}$ indicates a poor overall fit of the model to the data with little differentiability between models. For this reason, we require $7000<\chi^{2}<10000$ (effectively $\chi^{2}_{dof}\sim 1$) for objects with $\Delta\chi^{2}<100$. This requires a good model fit for objects with lower differentiability to be determined as a robust redshift.
[…]
We consider all BGS spectra to yield a robust redshift if they meet these conditions:
  1. 1.
    SPECTYPE = ‘GALAXY’
  2. 2.
    $z<1.5$
  3. 3.
    $\mathrm{SNR}_{z}>0.81$
  4. 4.
    $\chi^{2}\geq 7000$
  5. 5.
    Either 5a or 5b must be satisfied:
    1. (5a)
      $\Delta\chi^{2}\geq 100$
    2. (5b)
      $40\leq\Delta\chi^{2}<100$ and $\chi^{2}<10000$
Figure 4: Examples of two BGS spectra that pass (top) and fail (bottom) our quality cuts plotted using the Prospect visual inspection tool. The red line is the spectrum with $\sim$8000 flux density measurements across the entire wavelength range. The green line shows the uncertainty in the flux density measurement at each wavelength. Common emission lines (dashed purple) and absorption lines (dashed green) are shown. The top postage stamp is the color composite image of the object from Legacy Survey DR9 imaging. The lower panel is a magnified region around significant absorption/emission lines. For quality spectra (top) the absorption lines are clearly distinguishable. For spectra with poor redshift estimations the emission lines or absorption lines appear nonphysical indicating a poor model fit. The DESINAME of each object is found in the top left corner of each figure and both spectra were obtained in SV1 as part of the ‘bright’ program.
In Figure 4, we show an example of a BGS spectrum that passes the cut and one that does not. The BGS spectrum in the top panel meets the quality redshift criteria and indeed has very high SNR, a strong continuum, along with good $\chi^{2}$ and $\Delta\chi^{2}$ statistics. For the spectrum shown in the bottom panel that fails the cut, we can see a clear continuum with a model that does not match the absorption features.
A majority of BGS spectra (98.3%) pass the cuts defined in this section. Based on these classifications, there are 20,825 unique BGS targets spectroscopically confirmed as galaxies in the DESI-COSMOS field and 17,770 unique BGS targets spectroscopically confirmed as galaxies in the DESI-XMMLSS field. The redshift distribution of BGS galaxies with reliable redshift estimations in the DESI-COSMOS field is shown in the right panel of Figure 2. In comparing our robustness criteria to the ‘fiducial’ criteria in Hahn et al. (2023), we estimate the catastrophic failure rate by comparing the redshift estimates for objects with multiple observations. Using the fiducial criteria of ZWARN = 0 and $\Delta\chi^{2}>40$, the main BGS sample retains 98.9% of its objects and has an implied catastrophic failure rate of 1.4% (144/10064). Our robustness criteria retains 98.3% of objects, with an implied catastrophic failure rate of 0.72% (73/9974).
4.2 Luminous Red Galaxies
[…]
In Figure 5, the LRG spectra largely populate the same accepted region as the BGS program. The upper sequence in the left panel of Figure 5 is populated by LRGs that were observed in SV1 with longer exposure times than the LRGs in SV3 and MAIN. This leads to LRGs with higher SNR and therefore higher $\chi^{2}$ values. LRG spectra with robust redshifts meet the following condition:
  1. 1.
    SPECTYPE = ‘GALAXY’
  2. 2.
    $z<1.6$
  3. 3.
[…]
  4. 4.
    $\chi^{2}\geq 7000$
  5. 5.
    Either 5a or 5b must be satisfied:
    1. (5a)
      $\Delta\chi^{2}\geq 100$
    2. (5b)
      $40\leq\Delta\chi^{2}<100$ and $\chi^{2}<10000$
[…]
Figure 6: Examples of two LRG spectra that pass (top) and fail (bottom) our quality cuts. The structure and color scheme is identical to that in Figure 4. For quality spectra (top) the absorption lines, line Ca II H and K, are clearly distinguishable. For spectra with poor redshift estimations the emission lines appear nonphysical indicating a poor model fit. The DESINAME of each object is found in the top left corner of each figure and both spectra were obtained in SV1 as part of the ‘dark’ program.
In Figure 6 we show an example of an LRG spectrum that passes the cut and one that does not. The LRG spectrum that passes the cut indeed has very high SNR, a strong continuum break at 4000 Å  and clearly identified Ca II H and K features. The spectrum that fails the cut, was rejected due to a low $\Delta\chi^{2}$ indicating the fitted features are difficult to differentiate from similar features from fits at another redshift.
There are 11,192 LRG targets spectroscopically confirmed as galaxies in the DESI-COSMOS field and 9,448 LRG targets spectroscopically confirmed as galaxies in the DESI-XMMLSS field with robust redshifts. The redshift distribution of the LRG sample is shown in the right panel of Figure 5.
We can compare our robustness criteria to the ‘fiducial’ criteria in Lan et al. (2023) and estimate the catastrophic failure rate by comparing the redshift estimates for objects with multiple observations. Using the fiducial criteria of ZWARN = 0, $z<1.5$ and $\Delta\chi^{2}>15$, the main LRG sample retains 99.5% of its objects and has an implied catastrophic failure rate of 0.56% (30/5304). Our robustness criteria retains 97.9% of objects with an implied catastrophic failure rate of 0.29% (15/5179).
4.3 Emission Line Galaxies
ELG spectra typically exhibit low SNR in the continuum (median SNR per pixel <1<1) with strong emission lines, primarily the [OII] doublet. These [OII] $\lambda\lambda 3726,3279$ Å lines are sufficient for estimating a redshift even without informative signal in the rest of the spectrum as is often the case at high redshift ($z>1$). In other cases, particularly at lower redshift, the continuum and associated features can offer sufficient signal for robust classification, even when [OII] lines are weak. The combination of the SNR measurements of the [OII] doublet along with $\Delta\chi^{2}$ has been proven sufficient to reliably classify ELG spectra with quality redshifts (Raichoor et al., 2023), according to the following relation:
$\rm{log}_{10}(SNR_{\rm{[OII]}})>0.9-0.2\rm{log}_{10}(\Delta\chi^{2})$ (3)
Instead of using a custom routine within the main pipeline, we use the [OII] flux measurements provided by FastSpecFit (Moustakas et al., 2023). As with the BGS and LRG targets, we only include ELGs with a SPECTYPE = ‘GALAXY’ classification. We exclude all ELGs with $z>1.6$ as the [OII] doublet along with other distinguishable features are redshifted to the far red or out of the DESI wavelength coverage.
[…]
We demonstrate how the ELG sample populates the [OII] SNR vs $\Delta\chi^{2}$ parameter space in Figure 7, where the ELG sample populates a locus that increases in [OII] SNR with increasing $\Delta\chi^{2}$. The sliding [OII] SNR vs $\Delta\chi^{2}$ cut allows for robust redshift estimations for spectra with low $\Delta\chi^{2}$ as long as the the [OII] SNR is sufficiently high. A $\Delta\chi^{2}\geq 25$ threshold is conservatively set to ensure spectra with confident fits are not cut regardless of the [OII] SNR.
Reducing sky fiber contamination is important for the ELG sample as sky subtraction of inherently low SNR spectra can cause catastrophic failures. Using a representative sample of sky fibers, we examine what fraction of sky spectra pass a [OII] SNR vs $\Delta\chi^{2}$ cut in Figure 8. A SNR cut in the r-band ($\textrm{SNR}_{\rm{r}}$) was determined to balance a reduction in the sky fiber contamination with a desire to maintain high completeness in the ELG sample. We find that an additional SNR cut ($\textrm{SNR}_{\rm{r}}<0.22$), excludes 98.9% of sky fibers that pass the [OII] SNR vs $\Delta\chi^{2}$ criteria. In total 99.97% of sky fibers are excluded with all ELG quality cuts applied implying a 0.026% catastrophic failure rate for low SNR ELG spectra.
[…]
Spectra in the ELG sample that satisfy the following conditions yield a robust redshift:
  1. 1.
    SPECTYPE = ‘GALAXY’
  2. 2.
    $z<1.6$
  3. 3.
    Either 3a or 3b must be satisfied:
    1. (3a)
      $\Delta\chi^{2}\geq 25$
    2. (3b)
      $\log_{10}(\rm{SNR}_{\rm{[OII]}})>0.9-0.2\log_{10}(\Delta\chi^{2})$
  4. 4.
    $\rm{SNR}_{r}>0.22$
An example of an ELG spectrum that passes the quality cuts is shown in the top panel of Figure 9. It has very strong emission lines and a well resolved [OII] doublet. The spectrum in the bottom panel of Figure 9 fails the quality cuts and has very weak signal with almost no detectable signal in emission lines.
Figure 9: Examples of two ELG spectra that pass (top) and fail (bottom) our quality cuts. The structure and color scheme is identical to that in Figure 4. For quality spectra (top) the emission lines like the [OII] doublet are clearly distinguishable. For spectra with poor redshift estimations the emission lines appear nonphysical indicating a poor model fit. The DESINAME of each object is found in the top left corner of the figure and both were observed in SV1 as part of the ‘dark’ program.
With these cuts, 80.9% of the primary ELG sample is retained. In total, there are 27,475 ELG galaxies with reliable redshift estimates in the DESI-COSMOS field and 20,529 in the DESI-XMMLSS field. The redshift distribution of the ELG sample is shown in the bottom panel of Figure 7.
We can compare our robustness criteria to the ‘fiducial’ criteria in Raichoor et al. (2023) and estimate the catastrophic failure rate by comparing the redshift estimates for objects with multiple observations. The fiducial criteria identifies reliable redshift estimates by their [OII] emission and $\Delta\chi^{2}$. This criteria retains 89.0% of its objects and has an implied catastrophic failure rate of 1.35% (157/11630). Our robustness criteria, requires the object to be spectroscopically confirmed as a galaxy and limits the redshift to below the sensitivity of DESI. Our criteria retains 80.9% of objects with an implied catastrophic failure rate of 0.68% (71/10428).
4.4 Quasars
Quasars span a large redshift range, as shown in Figure 10. This expansive redshift range leads a diverse sample with various broad emission lines. This spectral diversity motivates the multi-step approach of using Redrock+afterburners in Section 3.1 to ensure quality spectral classification and redshift estimations for this sample. We apply MgII and QuasarNET afterburners to all objects targeted as a quasar or objects from other classes spectroscopically confirmed as a quasar (i.e., SPECTYPE = ‘QSO’).
[…]
Figure 11 shows an example of a red quasar that was incorrectly redshifted and misclassified as a galaxy by Redrock but was reclassified as a quasar with afterburners. A new redshift using Redrock was determined using the QuasarNET prior.
Figure 11: A quasar spectrum with a corrected redshift estimation and spectral classification after being processed through the MgII and QuasarNET afterburners. The original spectrum analysis from Redrock yielded a redshift estimation of z = 1.68 with SPECTYPE = ‘GALAXY’. With clear broad line emission, QuasarNET detected an incorrect classification of galaxy and the spectral classification was switched to quasar. Redrock also provided a more accurate redshift estimation of z = 1.4138 after using the QuasarNET classification and redshift as a prior. The DESINAME of this object is found in the top left corner of the figure. This object was observed in SV1 as part of the ‘dark’ program.
The primary quasar population for the DESI-COSMOS field consists of 3,086 quasars while the DESI-XMMLSS field consists of 3,507 unique quasars. Using visual inspection of repeat observations, the combination of Redrock with these two afterburners, we report a catastrophic failure rate of $\sim 2.6\%$.
4.5 Stars
For the stellar population in our catalogs, we report the radial velocity and physical stellar parameters given by RVSpecfit (Koposov et al., 2011; Koposov, 2019; Koposov et al., 2024a, 2025). RVSpecfit fits stellar templates directly at the pixel level in order to determine the objects radial velocities and other physical parameters. This direct-pixel level fitting allows for more accurate uncertainty estimates and can account for the diversity of stellar objects.
In order for an object to be fit by RVSpecfit, it must satisfy the following conditions:
  1. 1.
    has MWS_ANY targeting bit OR
  2. 2.
    has SCND_ANY bit set OR
  3. 3.
    classified as STAR by Redrock OR
  4. 4.
    has radial velocities within -1500 and 1500 km/s
  5. 5.
[…]
From the primary selections, we report the RVS measurements for 19,399 stars in the DESI-COSMOS field and 10,998 in the DESI-XMMLSS field.
4.6 Other Target Classes
The DESI-COSMOS and DESI-XMMLSS catalogs include smaller samples of other targets from secondary and special programs. The description of all these samples is given in Appendix A. With the framework established for LRGs, ELGs, and QSOs, quality cuts are applied to each sample depending on target class. Each sample was analyzed to determine appropriate quality cuts by examining the prominent features, redshift range, SNR per pixel, and spectroscopic classification. Each sample is then sorted into one of four groups for redshift quality assessment. For target classes with high SNR and strong continuum features, we impose quality cuts using $\Delta\chi^{2}$ and $\chi^{2}$ similar to the BGS and LRG samples. For targets that have low SNR but prominent features, an emission line quality cut is imposed as with the ELG sample. Secondary and special program spectra that have been targeted or spectroscopically confirmed as quasars are treated using the same afterburners as the primary quasar sample. Secondary and special program spectra that are stellar targets or spectroscopically confirmed as stars are processed through the RVS pipeline. A full list of all target classes with the number of robust redshifts for both fields can be found in Appendix B.
4.6.1 Secondary and Special Continuum-Characterized Samples
The secondary and special program samples described in this section have spectra that exhibit strong continuum and typically high SNR. Since these spectra have properties that are ‘LRG-like’, spectra in these samples have a robust redshift estimations if they meet these criteria:
  1. 1.
    SPECTYPE = ‘GALAXY’
  2. 2.
    $z<1.6$
  3. 3.
[…]
  4. 4.
    $\chi^{2}\geq 7000$
  5. 5.
    Either 5a or 5b must be satisfied:
    1. (5a)
      $\Delta\chi^{2}\geq 100$
    2. (5b)
      $40\leq\Delta\chi^{2}<100$ and $\chi^{2}<10000$
In order to demonstrate the robustness of applying these criteria to these samples, we show one larger subsample that falls into this continuum-characterized group.
The LOWZ sample consists of low redshift, dwarf galaxies intended to supplement the BGS sample at fainter magnitudes. This sample is intended to constrain the dwarf galaxy luminosity function and identify possible transient hosts.
The LOWZ sample has high SNR per pixel with a strong continuum similar to the LRG population, motivating a $\Delta\chi^{2}$ vs $\chi^{2}$ cut. As seen in Figure 12, a majority of the LOWZ sample falls within the accepted range of $\Delta\chi^{2}$ vs $\chi^{2}$. With all cuts imposed, 95.6% of LOWZ objects pass our robustness criteria.
Sample Quality Redshifts
PRINCIPAL (LRG) 30086
LOWZ_TIER3 1857
LOWZ_FAINT 1803
BGS_BGS_BRIGHT 1748
DESI_LRG 1079
LOW_Z 750
BGS_BGS_FAINT 665
LOWZ_TIER2 642
LOWZ_BRIGHT 569
BGS_BGS_FAINT_HIP 264
UNWISE_BLUE 185
LOWZ_TIER1 145
UNWISE_GREEN 38
Table 2: Secondary and Special Continuum Samples. The number indicates the number of robust redshifts in the DESI-COSMOS field. These samples observed in secondary and special programs exhibit spectral properties similar to the primary LRG and BGS sample. Redshift quality was assessed using similar algorithms to the LRG samples.
Figure 12: Distribution of the LOW_Z sample in the DESI-COSMOS field in $\chi^{2}$ vs $\Delta\chi^{2}$ parameter space. The red dashed line displays the $\chi^{2}$ and $\Delta\chi^{2}$ quality cuts. Spectra in the upper right are assumed to have quality redshifts.
A summary of all ‘Continuum-Characterized’ target classes is shown in in Table 2. Spectra in this group all are subject to the quality cuts outlined above.
4.6.2 Secondary and Special Emission Line Characterized Samples
These secondary and special program samples are characterized by prominent emission lines such as the [OII] doublet and [OIII]. Spectra in these target samples have redshift quality cuts modeled after the primary ELG sample and have quality redshifts if they meet these criteria:
  1. 1.
    SPECTYPE = ‘GALAXY’
  2. 2.
    $z<1.6$
  3. 3.
    Either 3a or 3b must be satisfied:
    1. (3a)
      $\Delta\chi^{2}\geq 25$
    2. (3b)
      $\log_{10}(\rm{SNR}_{\rm{[OII]}})>0.9-0.2\log_{10}(\Delta\chi^{2})$
  4. 4.
    $\rm{SNR}_{r}>0.22$
To demonstrate the robustness of applying these criteria to these samples, we examine two of the larger subsamples in the emission line group, SNe Hosts and the [OIII] sample for the DESI-COSMOS field.
The SNe Host program is designed to target the host galaxies of SNe Ia in order to obtain redshifts for dark energy constraints. Targets in these programs are similar to the ELG samples with prominent [OII] emission. In Figure 13, an example of a SNe host spectrum is shown with strong [OII] emission lines.
The [OIII] sample is comprised of two samples: low redshift galaxies selected using photometry from the Merian project (Luo et al., 2024) and a sample of galaxies intended for LSST-like weak lensing analysis (See Appendix A). The [OIII] emission lines for these samples provide a clean feature for quality redshifts.
Figure 13: Example of a spectrum in our emission line sample for secondary and special program targets. The structure and color scheme is identical to that in Figure 4. A representative spectrum for the secondary and special program targets that use the strength of the [OII] doublet for determining the redshift quality is shown. The DESINAME of this object is found in the top left corner of the figure. This spectrum was observed in SV1 as part of the ‘other’ program.
In Figure 14, spectra populate the accepted regions for both the [OII] and [OIII] samples, similar to the ELG sample. We summarize all the secondary and special programs that are included in each sample and report the number of quality redshifts for each sample in Table 3.
Figure 14: Statistics of the secondary and special emission line samples in the COSMOS field (See Table 3 for samples in the [OII] and [OIII] samples) Left: The distribution of the secondary SNe Ia Host sample in SNR[OII] vs $\Delta\chi^{2}$ parameter space. The dashed red lines indicate the cuts in SNR[OII] and $\Delta\chi^{2}$ where excluded spectra are located in the bottom left section. Right: The distribution of the secondary and special [OIII] sample in SNR[OIII] vs $\Delta\chi^{2}$ parameter space. The dashed red lines indicate the cuts in SNR[OIII] and $\Delta\chi^{2}$ where excluded spectra are located in the bottom left section.
Sample Quality Redshifts
[OII] Objects
ELG_PRINCIPAL 56346
ELG (SPECIAL) 13672
FILLER_HIP 3742
FILLER_LOP 1676
HSC_HIZ_SNE 590
DESI_ELG_LOP 539
DESI_ELG_HIP 411
MERIAN (z $\geq$ 0.5) 195
WISE_VAR_QSO (SPECTYPE = ‘GALAXY’ ) 184
PSF_OUT (SPECTYPE = ‘GALAXY’ ) 153
DESI_ELG 84
STRONG_LENS 17
[OIII] Objects
MERIAN (z << 0.5) 1180
DESI_DEEP_HIP 640
DESI_DEEP_LOP 455
DESI_DEEP_HIP2 118
Table 3: Secondary and Special Emission Line Samples. The number indicates the number of robust redshifts in the DESI-COSMOS field. These targets used [OII] or [OIII] emission line SNR to determine spectroscopic redshift quality.
4.7 Secondary and Special Quasar samples
Target classes in Table 4 are objects that were targeted as a quasar or quasar-like object and/or have been spectroscopically confirmed as quasars. All secondary and special program targets are processed through the MgII and QuasarNET afterburners to confirm spectral classification.
Sample Quality Redshifts
WISE_VAR_QSO 1215
PSF_OUT(]textttSPECTYPE=‘QSO’) 867
DESI_QSO 368
ISM_CGM_QGP 53
QSO_RED 10
QSO_Z5 2
LOW_MASS_AGN 1
Table 4: Secondary and Special Quasar Line Samples. The number indicates the number of objects with robust redshifts in the DESI-COSMOS field. Spectroscopic redshift quality was assessed using the same algorithms as the primary quasar sample.
4.8 Secondary and Special Stellar samples
Target classes in Table 5 are objects that were targeted as stars and/or have been spectroscopically confirmed as stars and processed through the RVSpecfit pipeline. A summary of all secondary and special program stellar targets is given in Table 5.
Sample Number
MWS_FAINT_BLUE 606
MWS_MWS_MAIN_BLUE 265
MWS_MWS_BROAD 191
MWS_FAINT_RED 64
MWS_MWS_MAIN_RED 64
MWS_CALIB 26
BACKUP_CALIB 26
FAINT_HPM 20
MWS_MWS_WD 16
MWS_MWS_BHB 7
MWS_RR_LYRAE 6
WD_BINARIES_BRIGHT 6
WD_BINARIES_DARK 6
HPM_SOUM 4
BHB 3
MWS_MWS_NEARBY 1
Table 5: Secondary and Special stellar samples. Similar to the primary stellar samples, objects in these samples are processed through the RVSpecfit pipeline.
5 Value-Added Catalog
We release all information in this study in two public catalogs, DESI-COSMOS and DESI-XMMLSS. These two catalogs can be found at the webpage for all DESI Value-Added Catalogs [^7] https://data.desi.lbl.gov/public/papers/c3/cosmos-xmmlss/. The data model for the catalogs can be found publicly here [^8] https://data.desi.lbl.gov/public/papers/c3/cosmos-xmmlss/README.md and all columns referenced in this section are described in Appendix C. For both fields, the data product is a FITS table with several extensions: SPECZ_CAT, DECALS_DR9_PHOT, DECAM_PHOT, HSC_WIDE_PHOT, and HSC_UD_PHOT. In the DESI-COSMOS catalog, we include two additional extensions, MERIAN_PHOT and COSMOS2020_PHOT.
The catalogs have two main components: the main spectroscopic redshift catalog and supplemental photometric measurements. The redshift catalogs consist of 311,102 unique objects which are row matched across the photometric extensions. Here we describe the data provided in the publicly available VACs for both the DESI-COSMOS and DESI-XMMLSS fields.
5.1 Redshift Catalog
The main data product of the spectroscopic catalog is found in the SPECZ_CAT extension. This extension models the DESI EDR spectroscopic redshift catalog data model [^9] https://desidatamodel.readthedocs.io/en/latest/index.html including target location, identification and targeting information. These columns include: DESINAME, TARGET_RA, TARGET_DEC, PMRA, PMDEC, OBJTYPE, DESI_TARGET, SCND_TARGET, MWS_TARGET, BGS_TARGET, SV1_DESI_TARGET, SV1_SCND_TARGET, SV1_MWS_ TARGET, SV1_BGS_TARGET, SV3_DESI_TARGET, SV3_SCND_TARGET, SV3_MWS_TARGET, SV3_BGS_TARGET, SURVEY, PROGRAM, CHI2, DELTACHI2, SPECTYPE, SUBTYPE, ZWARN, IS_QSO_MGII, IS_QSO_QN_NEW_RR, and TSNR2_BGS/LRG/ELG/QSO/LYA. Information indicating if the target has corresponding RVS measurements or photometry from the appended photometric catalogs is found in columns: HAS_RVS, HAS_DECALS_PHOT, HAS_DECAM_PHOT, HAS_HSC_UD_PHOT,HAS_HSC_ WIDE_PHOT, HAS_MERIAN_PHOT, HAS_COSMOS2020_PHOT.
5.1.1 Unique Targets
[…]
5.1.2 Targets observed in multiple programs
For objects observed in multiple epochs, we introduce several new parameters from the Redrock and quasar afterburners pipeline, attributed to the best available spectrum. Barring exceptions defined below, we defined the best available spectrum as the spectrum with the highest $\Delta\chi^{2}$ parameter. We merged redshift estimations, excluding spectra that do not have a quality redshift. For all redshifts, we report two quantities:
  • BEST_Z: The mean of all quality redshifts (QUALITY_Z = 1). For targets only observed in a single epoch, the BEST_Z is simply the redshift estimation from the data pipeline.
  • DZ: The largest difference in quality redshifts for the targets observed in multiple epochs, where $dz=\frac{z_{max}-z_{min}}{1+z_{min}}$. This parameter is set to 0 if the object was observed only in one epoch.
For all parameters associated with Redrock and quasar afterburners, we simply report the values associated with the best quality spectrum. The SURVEY and PROGRAM columns indicate the epoch from which the best spectrum was observed. We also include two columns describing the total coadded exposure time of the best spectrum, TOTAL_COADD_EXPTIME, and number of epochs of observation for the target, N_EPOCHS. All targeting information from each epoch will be included in the corresponding target mask columns. The reported emission line flux, emission line uncertainty and stellar fit parameters from the RVS pipeline represent the error weighted mean and error on the weighted mean over all quality spectra.
Visual inspection was performed on spectra with quality redshifts that disagree ($dz>0.0033$). For cases where only two observations were made of the same object, visual inspection was performed. If the visual inspection yielded a confident redshift (VI_quality $\geq 2.5$), the redshift from visual inspection is given in the BEST_Z column. If the visual inspection did not yield a confident redshift (VI_quality <2.5<2.5), the QUALITY_Z bit is flipped to False and the BEST_Z column reports the redshift from the spectrum with the highest $\Delta\chi^{2}$. In cases with more than two observations of the same object with only one discrepant redshift estimate, we simply took the redshift estimate identified by the majority of observations. In total, 70,799 objects had repeat observations with robust redshift classifications, for a total of 180,381 individual spectra. Of those individual spectra, the merging process revealed that 179,967 were given the correct redshift (99.77% purity) from the automated classification scheme.
A full list and description of each column can be found in Appendix C.
5.2 Photometry Extensions
In order to supplement the well-described target selection algorithms in Appendix A, we provide row-matched photometry from various imaging surveys for all targets. We matched all spectroscopic objects within $1^{\prime\prime}$ to the public imaging data sets, taking the closest match when multiple objects were located within $1^{\prime\prime}$ of the spectroscopic object.
The photometry includes DECaLs DR9101010https://www.legacysurvey.org/dr9/description/, DECaLS DR10111111https://www.legacysurvey.org/dr10/description/, Hyper Suprime-Cam PDR3 Ultra Deep/Wide (Aihara et al., 2022), photometry from the public COSMSOS2020 photometric catalogs (Weaver et al., 2022) and finally, photometry from the Merian medium-band survey (only for DESI-COSMOS) from the Victor M. Blanco telescope (Luo et al., 2024). The appended photometry is provided in extensions to the SPECZ_CAT for each photometric sample:
  • HDU2:DECaLS DR9 is an extension containing row matched magnitudes, flux estimates ($g,r,z,W1,W2$), flux uncertainties and selected columns from the public DECaLs DR9 catalog [^12] https://www.legacysurvey.org/dr9/files/#sweep-brickmin-brickmax-fits.
  • HDU3:DECaLs DR10 is an extension containing row matched magnitudes, flux ($g,r,i,z,W1,W2$), flux uncertainties and selected columns from the public DECaLs DR10 data [^13] https://www.legacysurvey.org/dr10/files/#sweep-brickmin-brickmax-fits.
  • HDU4:HSC PDR3 ULTRA DEEP is an extension containing row matched photometry from HSC Ultra Deep PDR3 and has selected data columns from the public data catalog [^14] https://hsc-release.mtk.nao.ac.jp/schema/#pdr3.pdr3_wide.forced2. The selected data columns cover photometry across ($g,r,z,i,y$) and include magnitudes, flux estimates, and flux uncertainties for different models.
  • HDU5:HSC PDR3 WIDE is an extension containing row matched photometry from HSC Wide PDR3 and has selected data columns from the public data catalog [^15] https://hsc-release.mtk.nao.ac.jp/schema/#pdr3.pdr3_dud_rev.forced2. The selected data columns cover photometry across ($g,r,z,i,y$) and include magnitudes, flux estimates, and flux uncertainties for different models.
  • HDU6:COSMOS2020 (DESI-COSMOS ONLY) is an extension containing row matched magnitudes, flux estimates, and flux uncertainties, photometric redshifts, galaxy parameter estimates from the COSMOS2020 public catalog including $\sim$30 broadband and medium-band optical and near-infrared filters (Weaver et al., 2022).
  • HDU7:MERIAN (DESI-COSMOS ONLY) is an extension containing row matched magnitudes, flux estimates, and flux uncertainties from the Merian Survey. The Merian filter set includes: N708 ($\lambda_{c}=7080\AA$ , $\Delta\lambda=275\AA$ ) and N540 ( $\lambda_{c}=5400\AA$ , $\Delta\lambda=210\AA$ ).
6 Example Applications
In this section, we provide example applications of the DESI-COSMOS and DESI-XMMLSS VAC catalogs demonstrating completeness with respect to the parent galaxy population, comparison to other spectroscopic compilations, photometric redshift quality, and galaxy cluster member identification.
6.1 Sample Completeness
A key benefit of adding high-fidelity spectroscopic information in the XMM-LSS and COSMOS fields is that both fields are observed by many different surveys, and so are used for cross-calibration and comparison across different data sets. For example low-noise photometric observations of objects in the deep drilling fields for DES (Childress et al., 2017) and LSST (Gris et al., 2024) provide an estimate of the true source fluxes. With an estimate of “truth”, these samples can be used to probe the selection function or can be used to infer the redshift distribution of samples used for cosmological analyses with photometric data (Kong et al., 2025).
Crucially, characterizing the selection function works well only for a high quality reference spanning a wide range in color-magnitude space. We opt to demonstrate the completeness of our DESI-COSMOS spectroscopic catalog using HSC Wide photometry, as this is most directly comparable to what will be available with Rubin.
We limit the completeness assessment to a radius 1.6 deg2 from the field center, RA, DEC = (150.1, 2.182). The full area of the DESI-COSMOS field includes more spectra, but with lower typical completeness. We apply a series of quality and footprint cuts on the HSC Wide sample photometry to obtain a “clean” sample based on pixel, centroiding, and flux flags, removing any objects that evaluate to TRUE for any of the following columns, for each filter in $\texttt{BAND}\in\{g,r,i,z,y\}$:
  • {BAND}_PIXELFLAGS_EDGE,
    {BAND}_PIXELFLAGS_BAD,
    {BAND}_PIXELFLAGS_INTERPOLATEDCENTER,
    {BAND}_PIXELFLAGS_SATURATEDCENTER,
    {BAND}_PIXELFLAGS_CRCENTER
  • {BAND}_SDSSCENTROID_FLAG
  • {BAND}_CMODEL_FLUX_FLAG,
    {BAND}_PSFFLUX_FLUX_FLAG,
    {BAND}_KRONFLUX_FLUX_FLAG,
    {BAND}_APERTUREFLUX_15_FLUX_FLAG
  • {BAND}_INPUTCOUNT_VALUE <3<3
Note that we do not apply a BRIGHTOBJECT_CENTER flag, as we find this to be too conservative a cut that removes 40% of the objects that have high-quality DESI redshifts. In total, these cuts remove $\sim 15\%$ of the objects in the DESI COSMOS region with $i<24.1$, mostly due to the pixel and centroiding flags.
We then characterize the fraction of all objects with clean photometry that have quality DESI redshifts in two different ways, emphasizing completeness in either apparent magnitude or apparent colors. Note that here we define completeness in the strict sense of what fraction of all objects in the imaging data have good DESI redshifts, which is distinct from and more stringent than the characterization of what parts of color-magnitude space have been observed with DESI. In reality, the latter is what has to be practically used for extrapolating existing spectroscopic catalogs outside of their domain of support, and a key challenge is identifying if and how such extrapolation is biased (McCullough et al., 2024; Weaverdyck et al., 2025).
Figure 15: Completeness of extended objects in DESI-COSMOS as a function of dereddened $i$ magnitude as measured from HSC. The gray curve gives the total number of extended HSC objects with clean photometry, the filled gray histogram gives the number of those objects with good DESI redshifts, and the blue curve gives the resultant fraction complete, i.e., the ratio of the two gray histograms.
Figure 15 shows the completeness as a function of dereddened $i$-band magnitude [^16] All HSC magnitudes are CMODEL magnitudes unless otherwise indicated., indicating that such samples are all but complete up to $i<17.5$, and with a majority of extended objects having spectroscopic classification up to $i<21.3$. The number of DESI spectra peaks at $i\sim 22.9$, and after this the completeness drops precipitously to effectively 0 at $i\sim 24.1$, as the number of photometric objects grows exponentially (shaded histogram).
This increase of galaxies at fainter magnitudes highlights a pernicious challenge for spectroscopic surveys: both the observational time required to get a good redshift per object and the number of objects increase geometrically as magnitude increases. The resulting lack of completeness at the faint end means that as photometric samples go deeper, the fractional number of spectra that are representative of such samples goes down rapidly. As a consequence, it is common to assume that various galaxy population characteristics (such as redshift) are dependent on color, but not on overall magnitude (Masters et al., 2015; Hildebrandt et al., 2020), such that observations from well-characterized bright populations can be applied to the faint objects that dominate photometric samples. However, it is known that such assumptions do not hold in practice (McCullough et al., 2024), such that representative spectroscopic measurements for faint samples are crucial for understanding their true properties.
Figure 16: Marginal density and completeness of clean HSC extended objects in the DESI-COSMOS area in the (gr,rz)(g-r,r-z) plane. The black curves give the $\{0.25,0.50,0.75,0.9,0.95,0.99\}$ object density iso-contours, while colors indicate the fraction of those objects with good DESI redshifts (limited to cells with at least 10 galaxies). The dashed purple line shows the boundaries of the main DESI ELG selection, for reference. Left: Sample selected with a limiting magnitude of $i<23$. Right: Sample selected with a limiting magnitude of $i<24.1$.
Figure  16 shows the DESI spectroscopic completeness of clean HSC extended objects projected into a simpler (gr,rz)(g-r,r-z) color space, and limited to $i<23$ magnitude (left) or $i<24.1$ magnitude (right). Contours show the density of the population, with colors illustrating completeness. There is clear inhomogeneity in the sampling of the grzgrz color space, with over-representation of some parts of the colorspace (e.g., due to the ELG selection, indicated via the purple dashed line).
Figure 17: Completeness properties in a 2D color manifold defined via grizy + $i_{\rm fiber}$ fluxes for a limiting magnitude of $i<24.1$ . Axes are arbitrary, but nearby SOM cells are also nearby in the SOM-learned colorspace. Left: Fraction of spectroscopically observed objects that pass our quality criteria ($\texttt{QUALITY\_Z}=True$) in each cell, identifying clear regions in color space with high failure rates and high redshift success rates. Center: Fraction of all clean HSC objects with good DESI spectra. Right: Average I_EXTENDEDNESS_VALUE of objects in each cell. DESI undersamples point sources, with relatively fewer good DESI spectra in cells with $\texttt{I\_EXTENDEDNESS\_VALUE}\sim 0$.
In Figure 17 we show completeness of a self-organizing map (SOM). This is a learned 2D manifold of the $grizy+i_{\rm{fiber}}$ HSC wide color space up to a limiting magnitude of $i<24.1$. We use the SOM algorithm described in Sánchez et al. (2020), which has several improvements to traditional SOMs, including a tunable hyperparameter that dictates how much the SOM cares about differences in the overall magnitude between objects vs. differences in the objects’ photometric colors, as well as accounting for measurement uncertainties in the fluxes.
The hyperparameter $\sigma_{F}$ defines a Gaussian prior on an overall amplitude rescaling of the flux in all bands, which is marginalized over when determining the “distance" between any given object and the SOM cells covering the magnitude-color-space. In the extremes, setting $\sigma_{F}=0$ allows for no rescaling, such that objects with the same color but rather different overall magnitudes will populate different SOM cells, whereas $\sigma_{F}\rightarrow\infty$ enables free rescaling of the overall flux without penalty, erasing magnitude information and resulting in a SOM that is only sensitive to color.
This avoids a common practical challenge when working with SOMs, which is whether to include only colors or also an overall magnitude when defining the relevant distance metric for learning the manifold.
Here we use $\sigma_{F}=0.4$, which corresponds to a $1\sigma$ penalty for an overall magnitude difference of 0.4. This gives a view that is more sensitive to colors than overall magnitude, and is largely complementary to the magnitude-only completeness characterization of Figure 15. This metric is more likely to group objects with like photo-z’s, which depend much more strongly on color than magnitude (c.f. Sánchez et al., 2020, for more details on the benefits of this SOM algorithm, particularly App. A).
Figure 17 illustrates the selection effects present in the DESI-COSMOS sample in the self-organizing map. The left plot shows clear regions with higher spectroscopic failure rates, requiring further spectroscopic follow-up. The center plot shows how quality DESI redshifts are highly concentrated in certain regions of color-space. The right plot shows how the SOM identifies and separates point sources naturally through the inclusion of both i and $i_{\rm{fiber}}$ fluxes. Cells characterizing point-like objects tend to be in regions with very low completeness, illustrating how DESI preferentially targets extended objects.
The central plot can be used to construct completeness weights (Lange et al., in prep). These would be computed as the inverse completeness fraction of each galaxy’s SOM cell, and enable one to reweight information coming from the DESI sample to better represent the respective imaging samples, accounting for selection along the HSC grizy and $i_{\rm{fiber}}$ axes. While this can remove some selection effects to make the DESI sample more representative of an LSST-Y1 like sample with $i<24.1$, it is only able to account for such effects that manifest in the this color-space, is limited by the resolution of the SOM, and does not properly account for objects observed with DESI for which we could not obtain a reliable redshift.
6.2 Spectroscopic Redshift Comparison
Complementing the DESI-COSMOS catalog, the spectroscopic redshift compilation from the COSMOS collaboration provides 48,575 quality redshifts ($CL~\textrm{(confidence level)}\geq 95\%$) from spectroscopy of 97,929 unique objects in the central 4 deg2 region (Khostovan et al., 2025). In the same footprint, the DESI-COSMOS catalog reports 86,049 quality redshifts. For objects classified as galaxies, the catastrophic rate ($\Delta v<1000$ km s-1) from the DESI catalog is expected to be less than 0.5%. Matching across the central 4 deg2 region, we find the 27,935 objects in common between the two catalogs. In Figure 18, a comparison between spectroscopic redshifts yields a 99.3% agreement within 1000 km s-1.
Figure 18: Spectroscopic redshift comparison between the DESI-COSMOS catalog and the zCOSMOS catalog(Khostovan et al., 2025). Top: The redshift distributions for the DESI-COSMOS and zCOSMOS catalogs. The DESI-COSMOS redshift distribution (black) encompasses the full 16 deg2 footprint. The DESI-COSMOS (blue) and zCOSMOS (red) show the redshift distribution for the central 4 deg2 covered by the zCOSMOS catalog. Bottom: The spectroscopic redshift comparison between the DESI-COSMOS and zCOSMOS catalogs. The comparison yields small bias ($b=2.01\times 10^{-4}$) and 99.3% of redshifts agree within 1000 km s-1.
The zCOSMOS catalog has 20,640 galaxies with high quality redshifts unique to the compilation, of which 3,333 are galaxies with $z>1.6$ that cannot be classified by the current version of the DESI pipeline due to limitations in the galaxy template set. The recent zCOSMOS catalog also includes physical parameter estimates such as stellar mass and star formation rates. The DESI-COSMOS offers 58,114 redshifts for $z<1.6$ galaxies in the central 4 deg2 footprint that are distinct from the COSMOS compilation. The DESI-COSMOS catalog includes an additional 131,702 unique redshifts in an extended 16 deg2. Both the DESI-COSMOS and DESI-XMMLSS catalogs provide prominent emission line flux estimates for all objects, well-defined target selection algorithms, and row matched photometry from public DECam, HSC and COSMOS2020 catalogs.
6.3 Photometric Redshift Comparison
The COSMOS2020 catalog (Weaver et al., 2022) provides a large collection of photometry within the 2 deg2 COSMOS field, completely within the 16 deg2 footprint of our catalog. Photometric redshifts were computed following the method in Laigle et al. (2016) where both galaxy and stellar templates are fitted to the observed photometry using the code LePhare (Arnouts & Ilbert, 2011; Ilbert et al., 2006). A comprehensive review of the photometric redshift (photo-z) measurements can be found in Weaver et al. (2022).
Using the matched COSMOS2020 catalog, we compared all robust spectroscopic redshifts to the measured lp_zBEST estimates of the photo-z (Figure 19). We only include spectra with extended morphologies for our calibration (SOLUTION_MODEL $\neq$ ‘PointSource’). For each object, we measure the $\frac{\Delta z}{1+z}$ between matched objects from the COSMOS2020 catalog using $z_{spec}=$ BEST_Z:
$\frac{\Delta z}{1+z}=\frac{z_{phot}-z_{spec}}{1+z_{spec}}$ (4)
We quantify the precision and accuracy of the photo-z measurements with our spectroscopic measurements through three metrics (dispersion, outlier fraction and bias). For dispersion, we use the normalized median absolute deviation (NMAD), defined as:
$\sigma_{\mathrm{NMAD}}=1.48\cdot\mathrm{median}\left(\left|\frac{\Delta z}{1+z}-\mathrm{median}\left(\frac{\Delta z}{1+z}\right)\right|\right)$ (5)
following Brammer et al. (2008) as it is less sensitive to outliers than the standard deviation or other similar measure of the dispersion. We report the fraction of outliers in two ways. First, denoted as $\eta$, the outlier fraction is defined as the fraction of galaxies whose photo-z deviate from their spec-z by $|\Delta z|>0.15\cdot(1+z_{spec})$ (Hildebrandt et al., 2012). Second, using $\sigma_{NMAD}$ as the characteristic scatter of the photo-z calibration, we determine the fraction of outliers that fall within $5\sigma_{NMAD}$. Lastly, the bias, $b$, is computed as the median of $\frac{\Delta z}{1+z}$.
Across the whole sample, as shown in Figure 19, we find $\eta=1.53\%$, $b=0.0004$ and $\sigma_{NMAD}=0.010$. In addition we find that 93.4% of photometric redshifts fall within $5\sigma_{NMAD}$ of the spectroscopic redshift. A summary of the comparison from the total sample and primary samples can be found in Table 6.
Figure 19: Comparison of DESI spectroscopic redshifts to photometric redshift estimates using COSMOS2020 LePhare photometric redshifts. Objects in the spectroscopic catalog have quality redshifts and are matched to the COSMOS2020 catalog within 1". The red dashed lines indicate the upper and lower $5\sigma_{NMAD}$ bounds. $\sim$93% of matched objects fall within the $5\sigma_{NMAD}$ bounds. The outlier fraction ($\eta$), bias ($b$), and $\sigma_{NMAD}$ are also given.
Sample Number $\sigma_{NMAD}$ $\eta$ $b$ $5\sigma_{NMAD}$
Total 18294 0.010 $1.53\%$ 0.0004 $93.7\%$
LRG 1135 0.008 $1.42\%$ 0.012 $76.5\%$
ELG 2615 0.004 $1.53\%$ 0.003 $85.9\%$
BGS 1422 0.010 $2.53\%$ -0.004 $78.1\%$
[…]
Examining the comparison of the full sample in Figure 19, there is a significant feature around spectroscopic redshift of z = 0.34, where there is notable disagreement between the redshift measurements. Inspecting a random sample of spectra from this ‘cloud’, the spectra exhibit strong Ca II H and K absorption and a strong continuum break around 5400 Å in the observer frame. An example spectrum is shown in Figure 20. As shown in the right hand panel of Figure 20, there is a gap in the intermediate band filters [^17] http://svo2.cab.inta-csic.es/theory/fps/index.php?mode=browse&gname=Subaru&gname2=Suprime&asttype= covering roughly 5400 Å to 5600 Å. When the Ca II H and K feature is located in this gap, the photometric redshift estimates tend towards higher redshifts that place the Ca II H and K break at the longer wavelengths in the gap.
Figure 20: Left: Spectrum of a galaxy with $z_{spec}=0.362$ and a photometric redshift estimate at $z_{phot}=0.4299$. The black dashed lines display strong Ca II H and K absorption is present indicating a confident spectroscopic redshift. The prominent absorption features are located around 5400 Å. The DESINAME of the public spectrum is located in the title of the left panel. The spectrum was observed during SV3 during the dark program. Right: Intermediate band filter response curves used for photometric redshift estimations from COSMOS2020 (Weaver et al., 2022). In both panels, the gap in intermediate band coverage is highlighted in yellow.
In addition to comparing photometric redshift estimates to objects with quality spectroscopic redshifts, we also examined the photometric redshift estimates of galaxies that have poor quality spectroscopic redshifts (QUALITY_Z = False). Figure 21 shows the comparison of the lp_zBEST estimates to those DESI spectroscopic redshifts. Not surprisingly, we find a twenty-fold increase in $\eta$, a large bias, and an intrinsic scatter that is forty-times larger than the case where DESI redshifts were reliable. Three distinct features are apparent in the comparison. First, a large fraction of photometric redshifts are above $z_{phot}>1.6$, which is beyond the range of DESI spectral templates for galaxies. It is likely that these photometric redshift estimates are accurate but DESI lacks the wavelength coverage to properly determine the redshift. Second, a significant number of galaxies have photometric redshift estimates consistent with zero or significantly below the DESI spectroscopic redshift estimate. Finally, a significant fraction of galaxies agree in the redshift estimates, implying the DESI redshifts are accurate even though we have low confidence in classification. It would be valuable to consider these objects that could not be classified by DESI in future photometric redshift training.
Figure 21: Photo-z comparison of objects with poor quality (QUALITY_Z = False )spectroscopic redshifts in the DESI-COSMOS catalog. The objects outside the red dashed lines deviate from the spectroscopic redshift by more than 15%. The top histogram shows the distribution of spectroscopic redshifts. The right side panel displays the distribution of photometric redshifts. The outlier fraction ($\eta$), bias ($b$), and $\sigma_{NMAD}$ are also given.
6.4 Galaxy Cluster Validation
In combination with galaxy cluster finding algorithms, galaxy redshifts can be used to spectroscopically confirm galaxy clusters, estimate cluster mass, and characterize member galaxies.
As an example, clusters in the SPIDERS cluster catalog (Kirkpatrick et al., 2021) were identified using CODEX (Finoguenov et al., 2020), an algorithm that identifies galaxy clusters from their X-Ray properties. The SPIDERS (Clerc et al., 2016) spectroscopic followup program was conducted within SDSS to obtain redshifts for galaxy cluster members to validate cluster membership and location, including observations in the XMM-LSS field. These SPIDERS observations were also used to estimate galaxy cluster mass (Farahi et al., 2016).
With a large spectroscopic sample in the XMM-LSS field, the DESI-XMMLSS catalog can be used to supplement the information about galaxy clusters in the SPIDERS cluster catalog and othe cluster catalogs, and potentially identify new clusters at higher redshift. To demonstrate the utility of the DESI-XMMLSS catalog, we selected one cluster with a high number of spectroscopically confirmed members from SDSS ($N_{spec}$ = 30) located in the XMM-LSS footprint. We report the number of DESI spectroscopically confirmed galaxies as function of redshift in the vicinity of the X-ray cluster center. Using the DESI-XMMLSS catalog, we clearly identify 35 galaxies within 5 arcmin as likely members of the cluster, with a mean redshift of $z=0.1398\pm 0.002$, as shown in Figure 22. Our estimate of the cluster’s redshift matches the spectroscopically validated redshift of the cluster at $z=0.1398\pm 0.0006$ from Kirkpatrick et al. (2021). Further studies into the redshifts of these cluster galaxies can inform our understanding galaxy cluster dynamics but are outside the scope of this paper.
Figure 22: The redshift distribution of galaxies as a function of radial separation from the center of SPIDERS Cluster 1_24090. We report the number of galaxies in redshift intervals of $\Delta z=0.01$, varying the radii from 1 (purple) to 5 (blue) arcmin around the cluster center at $\alpha=33.66^{\circ},\delta=-4.548^{\circ}$. A clear cluster is identified with a mean redshift of $z=0.1398\pm 0.002$, matching the spectroscopic redshift of the cluster from external spectroscopic catalogs.
7 Summary
This paper presents the DESI spectroscopic redshift catalogs for the DESI-COSMOS and DESI-XMMLSS fields. We developed robust spectroscopic redshift classification based on the spectroscopic properties of each target class in these two fields. In combination, the two catalogs present spectroscopic redshifts for more than 311,000 unique objects across two 16 deg2 fields with over 262,000 quality redshifts. The spectroscopic redshifts in this catalog are highly robust, yielding less than 1% catastrophic failure rates across various samples. These samples include, but are not limited to, blue-star-forming galaxies, passive red galaxies, active galactic nuclei, Type Ia supernovae hosts, low redshift dwarf galaxies and many other target selections. Lastly, we showed the utility of these heterogeneous samples through broad applications to completeness studies, spectroscopic and photometric redshift validation, and galaxy cluster confirmation.
As of the writing of this paper, DESI is progressing into its fourth year of observations, with regular data releases planned over the coming years. Future data from DESI releases will include spectra and redshifts for more galaxies, stars and quasars in these fields. In addition, the existing data on LAEs and LBGs will continue to be studied and characterized. The additional observations and LBG/LAE characterizations in these fields will be appended to these catalogs regularly as these data are processed and assessed for redshift robustness. The data model and redshift robustness procedures will minimally change as new data is added. In this way these catalogs will serve as a living record of DESI observations that will continue to serve as a resource for the DESI collaboration and the broader astronomical community.
ACKNOWLEDGEMENTS
The work of Joshua Ratajczak and Kyle Dawson was supported in part by U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award No. DESC0009959.
This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/501100011033), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies. The authors are honored to be permitted to conduct scientific research on I’oligam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.
DATA AVAILABILITY
All data points shown in the published graphs are available in the public VACs found here [^18] https://data.desi.lbl.gov/public/papers/c3/cosmos-xmmlss/. Spectra shown are available here [^19] https://data.desi.lbl.gov/public/dr1/spectro/redux/iron/healpix/ as part of the public DESI data release. DESINAME and observational program for each spectrum is included in the figure caption for traceability. Supplemental data for selected plots are publicly available at https://doi.org/10.5281/zenodo.16808719.
Appendix A Target Selections
The algorithms for targeting various classes of galaxy and quasar targets described in §2 and several additional classes of object for calibration are all recorded in the various DESI_TARGET bits. A brief description of those objects, taken directly from their associated targeting papers, is found below with a full description found in the cited papers.
PRIMARY TARGETS: The following targets are DESI’s primary targets with a brief description of their targeting algorithm outlined below:
  • LRG (Zhou et al., 2023): The LRG targets are selected using optical photometry in the grzgrz bands and near-infrared photometry in the WISE $W1$. The LRG selection cuts for the South are:
    • $z_{\rm{fiber}}<21.60$
    • $z-W1>0.8\cdot(r-z)-0.6$
    • gW1>2.9g-W1>2.9 OR rW1>1.8r-W1>1.8
    • $r-W1>1.8\cdot(W1-17.14)$ AND
    • rW1>W116.33)r-W1>W1-16.33) OR rW1>3.3r-W1>3.3
    where $g,r,z$, and $W1$ are magnitudes and $z$-fiber is the $z$-band fiber magnitude, i.e., the magnitude corresponding to the expected flux within a DESI fiber.
    The photometry in the North is slightly different from the South, and the selection cuts are tuned to match the number density and the redshift distribution in the South. The cuts for the North are:
    • $z_{\rm{fiber}}<21.61$
    • $(z-W1)>0.8\cdot(r-z)-0.6$
    • (gW1)>2.97(g-W1)>2.97 OR (rW1)>1.8(r-W1)>1.8
    • $(r-W1)>1.83\cdot(W1-17.13)$ AND
    • (rW1)>W116.31)(r-W1)>W1-16.31) OR (rW1)>3.4(r-W1)>3.4
  • ELG (Raichoor et al., 2023): The Main Survey ELG selection cuts are of three kinds: (1) quality cuts to ensure that the photometry is reliable; (2) a cut in the $g$-band fiber magnitude; (3) a selection box in the grg-r vs. rzr-z diagram. Likewise there is an extension of cuts for the ELG_VLO and ELG_LOP classes summarized below in Table 7.
    Sample Cut Comment
    ELG brick primary = True Brick primary selection
    $nobs_{grz}>0$ Observed in grzgrz-bands
    $flux_{grz}\cdot\sqrt{flux\_ivar_{grz}}>0$ Positive SNR in grzgrz-bands
    $(maskbits\&2^{1})=0$, $(maskbits\&2^{12})=0$, $(maskbits\&2^{13})=0$ Not close to bright star/galaxy
    ELG_LOP $(g>20)\ and\ (g_{fib}<24.1)$ Magnitude constraints
    0.15<rz0.15<r-z Color condition 1
    $(g-r)<0.5\cdot(r-z)+0.1$ Color slope condition 1
    $(g-r)<-1.2\cdot(r-z)+1.3$ Color slope condition 2
    ELG_VLO $(g>20)\ and\ (g_{fib}<24.1)$ Magnitude constraints
    0.15<(rz)0.15<(r-z) Color condition 1
    $(g-r)<0.5\cdot(r-z)+0.1$ Color slope condition 1
    $(g-r>-1.2\cdot(r-z)+1.3)\ and\ (g-r<-1.2\cdot(r-z)+1.6)$ Color slope condition 2
    Table 7: The cuts are the same for the North and South-DECaLS/DES regions. We use the following definitions: ${grz}=22.5-2.5\cdot log10($flux_grz/mw_transmission_grz), $g_{fib}=22.5-2.5\cdot log10($fiberflux_g/mw_transmission_g). The brick primary, nobs_grz, flux_grz, fiberflux_g, flux_grz, mw_transmission_grz, maskbits columns are described here: https://www.legacysurvey.org/dr9/catalogs/
    ELG_HIP, is a ten percent random subsampling of the ELG_LOP and ELG_VLO samples, with the same fiber assignment priority as one of the LRG targets
  • QSO (Chaussidon et al., 2023):The DESI quasar target selection is a combination of optical-only and optical+IR colors. Two colors are used, grzWgrz-W vs. gzg-z where grzgrz is a weighted average of the grzgrz band fluxes and W a weighted average of $W1$ and $W2$ fluxes:
    $\texttt{flux}_{grz}=[\texttt{flux}_{g}+0.8\cdot\texttt{flux}_{r}+0.5\cdot\texttt{flux}_{z}]/2.3$ (A1)
    $\texttt{flux}_{W}=0.75\cdot\texttt{flux}_{W1}+0.25\cdot\texttt{flux}_{W2}$ (A2)
    To improve the success rate for DESI, we use a machine-learning algorithm based on Random Forests (RF).we restrict the selection to objects with stellar morphology (’PSF’ in DR9), to avoid an almost 10-fold contamination by galaxies that otherwise enter our selection region, and we impose $16.5<rAB<23.0$. In addition, to reject stars, we apply a cut on the (WISE) magnitudes ($W1<22.3$ and $W2<22.3$). This cut is particularly efficient at getting rid of stars in the Sagittarius Stream, a region which exhibits an overdensity of QSO targets. We also require that the targets are not in the vicinity of bright stars, globular clusters, or large galaxies. Such “masked” sources have MASKBITS of 1, 12 or 13 set in Legacy Surveys catalogs.
  • BGS (Hahn et al., 2023):
    • BGS_ANY: The BGS_ANY sample is any object selected from the BGS criterion outlined below. First a spatial mask is applied. masks are compiled using 773,673 Gaia DR2 objects with $G_{\rm{Gaia}}<13$ and 3,349 Tycho-2 objects with visual magnitude brighter than MAG_VT << 13. The masks have radii:
      $R_{BS}(m)=815\cdot 1.396^{-m}arcsec$ (A3)
      where m is either $G_{\rm{Gaia}}$ or Tycho-2 MAG_VT magnitude. We use $G_{\rm{Gaia}}$ when both are available.We do not apply spatial masking around stars fainter than 13th magnitude. Next, for GCs, we apply a circular mask with radius defined by the major axis of the object. We apply this masking around all of the GCs and PNe in the list compiled from the OpenNGC catalog. In total, our bright star and GC masks exclude 0.87 and 0.01% of the initial area, respectively.
      For the BGS targets, we only want to include galaxies and exclude stars. An object is considered a BGS target if either of the following conditions are met:
      1. 1.
        object is not in the Gaia catalog;
      2. 2.
        object is in Gaia and has $(G_{\rm{Gaia}}-r_{\rm{raw}})>0.6$.
      $G_{\rm{Gaia}}$ is the G band magnitude from Gaia and $r_{\rm{raw}}$ is the LS r band magnitude that is not corrected for galactic extinction. Our ($G_{\rm{Gaia}}-r_{\rm{raw}}$) criterion takes advantages of the fact that $G_{\rm{Gaia}}$ is measured assuming that the object is a point source for a (narrow) diffraction-limited point spread function (PSF) measured from space.
      Some of the objects in the LS are imaging artifacts or fragments of ‘shredded’ galaxies. In order to remove these spurious objects from our BGS target catalog, we apply the fiber magnitude cut (FMC) below:
[…]
      $r_{\rm{fiber}}$ is the $r$-band fiber magnitude derived from the predicted $r$-band flux of the object within a 1.500 diameter fiber and $r$ is the total $r$-band magnitude
      We want BGS Bright to be complete in all three optical bands of its imaging: $g$, $r$, and $z$. We therefore require that there is at least one photometric observations in each of the bands:
      $nobs_{i}>0\ for\ i\ =g,r,z$ (A5)
      $nobs_{i}$ represents the number of observations (images) at the central pixel of the source in each band. We also exclude spurious objects (e.g. imaging artifacts or stars) with extreme colors by requiring:
      1<(gr)<4-1<(g-r)<4 (A6)
      1<(rz)<4-1<(r-z)<4 (A7)
      Lastly, flux from very bright objects on neighboring fibers can contaminate the traces of faint objects on the spectrograph CCD, and pollute their observed fluxes. This is particularly problematic at >9000>9000 Å, where 10% of the flux of the contaminating source is scattered into the wings of its PSF. Since the faintest fiber magnitudes in BGS have $r_{\rm{fiber}}\approx 21.5$, we remove all objects that meet
      $(r>12)\ \&\ (r_{\rm{fibertot}}<15)$ (A8)
      from the BGS target catalog. Here, $r_{fibertot}$ is the total fiber magnitude derived from the predicted $r$-band flux within a 1.500diameter fiber from all sources.
    • BGS_ BRIGHT: For BGS_BRIGHT targets we impose a simple $r<19.5$ magnitude limit to select the BGS Bright targets.
    • BGS_FAINT: BGS includes galaxies with magnitudes fainter than $r>19.5$. This fainter sample will substantially increase the overall BGS target density and, thus, enable small-scale clustering measurements with higher signal-to-noise. The selection for these is following $r_{\rm{fiber}}$-color cut:
[…]
      Here, $W1$ is the magnitude in the WISE $W1$ band, 3.4µm at 6.100 angular resolution. We also include a $19.5<r<20.175$ magnitude limit in order to satisfy a $\sim 1,400$ targets/deg2 constraint on the total target density imposed by the survey fiber budget allocated to BGS.
    • BGS_WISE: To select targets with AGN, we exploit optical and infrared colors that trace the signatures of hot, AGN-heated dust in the spectral energy distribution. The primary AGN selection criteria are:
        (zW2)(gr)>0.5(z-W2)-(g-r)>-0.5
        (zW1)(gr)>0.7(z-W1)-(g-r)>-0.7
        W1W2)>0.2W1-W2)>-0.2
        $(G_{\rm{Gaia}}-r)<0.6$
      We also require a SNR >10>10 detection in both $W1$ and $W2$ bands to ensure a robust constraint in the infrared regime, and apply quality and magnitude cuts.
    • BGS_FAINT_HIP: In the DESI catalogs these BGS_FAINT targets with higher priority are labeled under the BGS_FAINT_HIP bitname.
  • MWS (Cooper et al., 2023):
    • MWS_ ANY: The MWS main sample is selected from the LS DR9 source catalog combined with astrometric measurements from Gaia EDR3. The following definition encompases any target with the MWS_ANY bit:
        $16<r<19$
        $r_{\rm{obs}}<20$
        type = PSF
        gaia_astrometric_excess_noise <3<3
        gaia_duplicated_source = False
        brick_primary = True
[…]
        {g,r}_flux >0>0
[…]
      The observed $r$-band magnitude is obtained from the LS flux in nanomaggies as $r_{\rm{obs}}=22.5-2.5\log_{10}(r_{flux})$. The extinction-corrected magnitude is computed as $r=r_{\rm{obs}}+2.5log10$(mws_transmission_r)
    • MWS_MAIN_BLUE The MAIN_BLUE subsample is approximately magnitude-limited selection, comprised mainly of main-sequence turnoff stars and bluer subgiants and defined by the color criterion:
      gr<0.7g-r<0.7 (A10)
      with $g$ defined in the same way as $r$ above.
    • MWS_MAIN_RED The MWS_MAIN_RED subsample is defined by the following additional criteria to the MAIN_ANY criteria :
        $g-r\geq 0.7$
        astrometric_params_solved $\geq 31$
        Gaia parallax $\pi<3\sigma_{\pi}+0.3$ mas
        Gaia proper-motion: $|\mu|<5\sqrt{f_{r}/f_{19}}$ mas/yr  $(f_{r}=10^{(}(22.5-r)/2.5)$
    • MWS_MAIN_BROAD
      The MWS_MAIN_BROAD sample comprises stars with color $g-r\geq 0.7$ that satisfy the same magnitude and data quality requirements as the other two categories, but do not satisfy one or more of the astrometric criteria that define main-red (stars that do not have well-measured astrometric parameters in the Gaia catalog are therefore included in this selection).
    • MWS_WD
      The MWS_WD has an extensive selection summarized here. White dwarfs are selected using a set of BP - RP color and $G$-band absolute magnitude criteria based on Gentile Fusillo et al. (2019) that identify the white dwarf cooling sequence in the Gaia photometric catalog alone. This selection is applied to the LS catalog using the properties of cross-matched Gaia EDR3 sources; the LS photometry is not used in this selection. Photometric measurements are taken from Gaia DR2 and astrometric measurements from EDR3.
        $G_{abs}>5$
        BPRP<1.7BP-RP<1.7
        $G_{abs}>5.93+5.047(BP-RP)$
        $G_{abs}>6(BP-RP)^{3}-21.77(BP-RP)^{2}+27.91(BP-RP)+0.897$
      These criteria are applied to a sample defined by extinction-corrected Gaia G flux (gaia_phot_g_mean_mag in the LS catalog) at high latitude, with mild parallax and proper-motion cuts to select against nearby luminous blue stars (early-type main-sequence stars, horizontal branch stars, and subdwarfs) as well as QSOs:
        $G<20.0$
        $|b|>20$ degrees
        $\pi/\sigma_{\pi}>1$ mas
        astrometric_params_solved $\geq 31$
        $|\mu|>2$ mas/yr
      We impose the following photometric quality cut, because failing this criterion results in poor astrometry:
        phot_bp_rp_excess_factor $1.7+0.06(BP-RP)^{2}$
      However, we retain objects that have reliable parallaxes and significant proper-motions that meet either of the following criteria:
        astrometric_sigma5d_max <1.5<1.5
        (astrometric_excess_noise $1)\ \&\ (\pi/\sigma_{\pi}>4)\ \&\ (|\mu|>10$ mas/yr)
    • MWS_RR_LYRAE
      We target Gaia DR2 sources with magnitudes $14<G<19$ that are classified as RR Lyrae by the general variability classifier and the Special Object Studies classifier (Holl, B. et al., 2018; Clementini, G. et al., 2019), comprising all objects from the table vari_rrylrae and any sources from the table vari_classifier_result for which best_class_name includes ”RR”.
    • MWS_NEARBY For MWS_NEARBY targets, we select dwarf stars within 100 pc of the Sun based only on apparent Gaia magnitude and parallax, allowing for moderate parallax uncertainties:
        $16<G<20$
        $\pi+\sigma_{\pi}>10mas$
        astrometric_params_solved $\geq 31$
      MWS_NEARBY targets are prioritized over all other MWS targets except MWS_WD and MWS_RR_LYRAE.
    • MWS_BHB BHBs are selected starting from the basic definition of main sample stars given at the start of this section (common to main-blue and main-red), with the following additional criteria:
        $G>10$
        $\pi\leq 0.1+3\sigma_{\pi}mas$
        $0.35\leq g-r\leq-0.02$
        $0.05\leq X_{BHB}\leq 0.05$
        $r-2.3(g-r)-W1_{faint}<-1.5$
      These criteria exclude nearby stars and select around the BHB locus in a combined LS and WISE color space, defined by:
      $X_{BHB}=(g-z)-[1.07163(g-r)^{5}-1.42272(g-r)^{4}\\ +0.69476(g-r)^{3}-0.12911(g-r)^{2}+0.66993(g-r)-0.11368]$ (A11)
      and
      $W1_{faint}=22.5-2.5\log_{10}(W1-3\sigma_{W1})$ (A12)
      where W1 and $\sigma_{W1}$ are the WISE 3.4µm flux and its error, respectively.
SECONDARY TARGETS: The following targets are DESI’s secondary targets with a brief description of their targeting algorithm outlined below. These targets are flagged as SCND_ANY bit in the main desi mask and can be found in the scnd mask:
  • UDG: Targets are selected from the SMUDGes catalog of the DECaLS data. SMUDGes analyzes the DECaLS data images to identify low surface, diffuse galaxies (central surface brightness >> 24.0 mag/sq. arcsec and half light radii >> 5.3 arcsec). The procedure involves various steps described in Zaritsky et al. 2019 and in another paper in progress. The specific subsample presented here are those ultradiffuse galaxies in the SMUDGes catalog that have gr<0.3g-r<0.3, making it more likely that they have emission lines from which measuring the redshift would be simplified. Some of these sources are already included in the BGS Bright and Faint samples, but many are not (we set override = 0 so as to not affect targeting for those galaxies previously identified).
  • QSO_RED: A selection of dust-reddened quasars with selection using DR8 quasars as a base sample:
    • $r$-band magnitude cut ($r<23$): ensures high SN sources observed
    • require MORPHTYPE==‘PSF’: to select QSO-like sources
    • remove sources within the main DESI survey color selection (0.4<rz<1.1&gr<1.3-0.4<r-z<1.1\ \&\ g-r<1.3): this reduces the chance of re-observing a blue QSO that will already be targeted in the main survey
    • require W1, W2 and W3 detections and a $SNR>3$ in all three bands
    • remove sources that do not satisfy the Mateos et al. (2012) WISE AGN wedge using the W1 - W2 and W2 - W3 colours: this important step significantly reduces the number of non-QSO contaminants
  • FIRST_MALS:These targets are faint radio sources that have 21-cm and OH absorption spectra from MALS generated by cross-matching the radio source positions with DR8 source locations with a positional accuracy of 1 arcsecond. Initially we are focusing on MALS points with RA in the range 120210120-210 deg and DEC in the range 10-10 to +20+20 deg. There are 85 MALS pointing overlapping with DESI foot prints in the above mentioned RA range. Within the field of view of MALS observations we have identified 19107 radio sources with peak flux density of >10>10 mJy using FIRST survey. The flux limit is set by the SNR requirements from the 21-cm absorption line searches. While cross-matching with DR8 catalog of DECALS we have identified 8300 sources that are listed in the above file. We just want to mention that about 10% of our MALS FOV have poor coverage in the FIRST survey.
  • MWS_CLUS_GAL_DEEP: We have 3 different classes of objects here, dwarf galaxies, globular clusters and open clusters. Since they have different properties, we select the targets with different criteria, detailed below
    • Dwarf Galaxies (DC):
      Two dwarf galaxies are selected for SV, Draco and Ursa Major II (Uma II).
      We select field centers to be $(ra+0,dec+0.4)$ where ra/dec are the center of the galaxies. This offset is chosen so that the dwarf center fall on one of DESI petal instead in the center of the field (to avoid possible fiber gaps). We first select targets from DECaLS DR8.Then targets are selected with 1.7 deg from the field center, with brick_primary = True and type=‘PSF’. The later is to ensure the targets are stellar like object. All photometry used for selection are dereddened using SFD98. We first select targets have $r>16$ and gr<1.2g-r<1.2.We then select targets with two passes:
      First pass, for stars that have a proper motion (pm) and parallax (w) measurements, we select targets with:
        Proper motion cut: $|\mu-\mu_{0}|<2$ where $\mu_{0}$ is the proper motion of the cluster
        Parallax cut: $\pi-3\cdot\sigma_{\pi}<1/distance$
[…]
      Second pass, for stars that have no PM or parallax measurements, we select targets with DECaLS photometry. Specifically we select targets using a metal poor isochrone to select stars are consistent to isochrone within certain width. Finally we combine the two passes and put stars with $G>18$ as high priority targets and $G<18$ as low priority targets. (Because $G<18$ stars should mostly observed by regular MWS SV). We have a total of 4079 targets for Draco and 2306 targets for Uma II.
    • Globular Clusters(GC):
      Two GC are selected, Pal 13 and M 53. M 53 field has two GC, M 53 and NGC 5053, which can be observed together with one DESI field.
      Selection for GC is very similar to DG, with the only exception is:
      1. 1.
        In the photometry cut, we used $r>16$ and gr<1.1g-r<1.1
      2. 2.
        In first pass, proper motion cut is $|\mu-\mu_{0}|<3$
      We have a total of 2741 targets in Pal 13 and 11994 targets in M53. Note that although there are a lot of targets in this field, if we remove the crowded cluster core, the target number drops to 4027 in total. Therefore, the outskirt of the GCs will still be able to observed by DESI.
    • Open clusters (OC): Fields centered around the central cluster coordinates using DECALS (and the Gaia astrometry/photometry in the DECALS files).
      For each cluster, two catalogs are created. One with high priority containing astrometric cluster members, and a second one with targets that match the g-r and r-z colors as a function of g magnitude.
  • LOW_MASS_AGN: We propose to target AGN candidates selected from the DESI tractor photometry in the z and WISE bands, in order to search for AGN in low-mass galaxies. The target selection steps are as follows:
    Consider sources with the following conditions:
    • brick_primary = 1, maskbits = 0, wisemask_w1 = 0 and wisemask_w2 = 0, anymask_g = 0 and anymask_r = 0 and anymask_z = 0
    • 0.02 $\leq$ z_phot_median $\leq$ 0.3
    • flux_X and flux_ivar_X columns selected where X = g,r,z,w1,w2,w3,w4
[…]
    • Removing extreme color sources. Keeping only sources with:
        $2.0\leq z-W1\leq 2.0$
        $2.0\leq W1-W2\leq 2.0$
        $2.0\leq W2-W3\leq 4.0$
    • Computing the absolute magnitude in $z$-band, MzMz from zzmag and z_phot_median or selecting low-mass faint galaxies, we make a cut in $Mz$ and $r$:
        $Mz\geq-21$
        $r\geq 19$
    • Using the new WISE color-color AGN diagnostics, we make two agn candidate selections:
        AGN1  selection: $(W1-W2<(0.5667\cdot(z-W1))+0.5)$ and $(W1-W2>(-1.5714\cdot(z-W1))+0.5)$ and $(W1-W2>(0.4375\cdot(z-W1))-0.90625)$
        AGN2  selection: (W1W2>0.4)(W1-W2>-0.4) and $(W1-W2>(0.2\cdot(W2-W3))-0.6)$ and $(W1-W2>(4\cdot(W2-W3))-8.2)$
    • We define three priority classes:
        High: Candidates that satisfy both agn1 and agn2 selection and have $W3\leq 17$.
        Medium: Candidates that satisfy both agn1 and agn2 selection and have $W3>17$
        Low: Candidates that satisfy only of the selections, i.e., agn1 or agn2, but not both.
  • FAINT_HPM:
    1. 1.
      HPM-G : All Gaia DR2 stars with $G\geq 19$ and Gaia proper motion >> 100mas/yr
    2. 2.
      HPM-N : Sources from the NOIRLab Science Catalog DR2 with NSC proper motion >> 100mas/yr and $(15<g<23)$ or $(15<r<23)$ or $(15<z<22))$ where g,r,z are NSC mags.
    3. 3.
      HPM-BD All brown dwarf candidates from WISE, DES, Ultracool, etc. (no proper motion cut)
  • LOW_Z_TIER_1/2/3: This program uses spare fibers in dark time to pursue a complete survey of the low-z Universe. After generating the list of objects passing the SAGA II cuts (Mao et al., 2021), the target lists are then sorted by the following three tiers:
    Tier 1: CNN predicted objects from SAGA II cuts + SAGA II extended cuts (10% extra in color and magnitude)
    • Sorting: pCNN value (CNN probability)
    • Density: 16 per square degree
    Tier 2: SAGA II cuts outside of BGS sample (r<19.5and 19.5<r<20.3r<19.5\ and\ 19.5<r<20.3 + color cuts)
    • Sorting: pSAT value (satellite probability from Mao et al. (2021))
    • Density: 123 per square degree
    • ($r_{fib}>23$): 3 per square degree
    Tier 3: BGS sample + SAGA II extended cuts
    • Sorting: random
    • Density: 200 per square degree
    • BGS galaxies: 72 per square degree
    • SAGA II extended cuts: 128 per square degre
  • BHB: A complement to the DESI Milky Way (MW) Survey by observing faint Blue Horizontal Branch (BHB) and RR Lyrae stars in the dark observing time of DESI (Byström et al., 2024). The target selection is as follows:
    1. 1.
      Selection of blue sources:0.3<gr<0-0.3<g-r<0
    2. 2.
      Magnitude limit: $19<g<21.5$
    3. 3.
      Color selection to separate BHB from BS/WD and QSO Poly1(gr)<rz<Poly2(gr)Poly1(g-r)<r-z<Poly2(g-r) where Poly1 and Poly2 are quadratic polynomials
    4. 4.
      WISE selection: $f_{W1}/f_{G}<0.3(g-r)+0.15+\sigma[_{W1}/f_{G}]$ where $f_{W1}$ and $f_{G}$ are fluxes in Gaia G and WISE W1 and $\sigma[_{W1}/f_{G}]$ is the uncertainty on the flux ratio. This is necessary to remove residual contamination from QSOs.
    5. 5.
      Quality cuts: ANYMASK G/R/Z and GALAXY maskbits are not set.
    As RR Lyrae are essentially pulsating counterparts to blue horizontal branch stars we also include a small number of distant RR Lyrae stars that have been identified by Gaia (Gaia Collaboration et al., 2019) and PS1 (Sesar et al., 2017) within the same magnitude limits.
  • PSF_OUT_BRIGHT/DARK: Program to obtain spectroscopy of all point sources of unusual color in order to find missing quasars and to be a discovery engine for new, interesting objects. We simply target any sources of type PSF that are inconsistent with the colors of main sequence stars and that are missed by other selection criteria. This includes the 18 deg2 at $15<r<19$ as bright time spare fibers, and the 130 deg2 at $16<r<22$ as dark time spare fibers.
  • HPM_SOUM:A selection of high proper motion stars. To make our target list, we used the list of HPM objects from Segev & Ofek (2019). We calculated the current locations of all the objects in the list (using the proper motion provided). We then cross-matched the current locations with the DR8 bricks corners to make a list of targets that are currently in the DR8 footprint.
  • HSC_HIZ_SNE: Our candidates are selected by the presence of a SN in the host galaxy within the HSC COSMOS deep field (Yasuda et al., 2019). There is no magnitude or color cut on the targeted host but we cut at $r=25$ for the host galaxies of less certain SNe. Internally the targets are prioritized according to their likely scientific value for constructing a Hubble diagram and measuring the dark energy equation of state. Highest priority went to host galaxies of SNeÍa observed with HST as part of programs GO14808 and GO15363 (PI Suzuki); priority then decreased in the order of: other likely SNe Ia; including the HST SNe Ia), possible SNe Ia, and less certain SNe Ia. We then removed the targets whose redshifts we had previously obtained with AAOmega (Sharp et al., 2006), and targets with redshifts from the literature. To make a more complete Hubble diagram, we also include host galaxies for confirmed orsuspected SNe Ia lacking redshifts with the SNLS D2 deep (Astier et al., 2006).
  • SN_HOSTS: A set of targets for SNe host to understand SN/host correlations, SNe Ia cosmology, fundamental plane astrophysics and other scientific pruposes. The selection is summarized below:
    1. 1.
      We computed a list of transients based on spectroscopically classified transients from,ZTF, PTF/iPTF, SDSS II, Greg Aldering’s sample including historical SNe, ZTF QSOs and made sure we have no duplicates
    2. 2.
      For the non-QSOs sources, we ran a host identifier code which: (a)looked at possible hosts from the local universe and (b) looked at possible hosts from dr8_north and dr8_south. "Possible hosts" are selected using two criteria: (1)light radius proximity (2)redshift proximity. When the two criteria do not agree, the host is ambiguous and we request to observe both possible hosts.
    3. 3.
      For the QSOs sources, we simply give their RA and DEC
    4. 4.
      We add 259 hosts from Greg Aldering’s sample which disagree with our host identifications, for reasons we believe are related to DR8 and should disappear in DR9.
  • STRONG_LENS:Our targets are selected from strong lensing candidate systems discovered in DESI Legacy Imaging Surveys. They target the $\sim 70$ lenses not currently in the DESI target list and the brightest lensed image for every source (typically one per lensing system) to the same depth as other DESI targets (4 passes, $NUMOBS=1$) in dark time, to be revisited based upon SV observations. We will provide coordinates for all our targets.
  • GAL_CLUS_SAT This sample is a volume limited complete spectroscopic sample of BCGs and the bright cluster members. The BCG is the SDSS-redMaPPer most probable central. To determine the fraction of SDSS-redMaPPer BCGs and members fibered by DESI, we do the following. We note that galaxies not fibered likely result from fiber collisions.
    • We identify SDSS-redMaPPer BCGs and galaxies with $P_{mem}>0.90$ in the DESI northern footprint.
    • We use fiber assignments provided to BGS and C3 by Jaime Forero-Romero todetermine which galaxies will be fibered. In that simulation all tiles for the entire surveyare fed into fiberassign in a single run (i.e., we do not simulate the survey progress nor any type of cadence to update the Merged Target List (MTL)). We use DR9 targets and incorporate multiple passes. Since some galaxies may be observed during dark time, we also determine which DESI LRG targets correspond to our galaxies.
    • For the BCGs we extend our analysis out to $z<0.35$. For the satellite galaxies, we only go out to $z<0.30$.
    • In total, there are 202 BCGs, 260 second brightest galaxies, and 10787 satellite galaxies with $P_{mem}>0.90$.
    This sample is all the satellite galaxies (excluding the 2nd brightest) in the galaxy cluster.
  • MWS_FAINT_BLUE/RED
    The faint-blue and faint-red samples extend the main-blue and main-red classes, respectively, to the fainter magnitude range $19<r<20$. The selection criteria are otherwise identical to those for the corresponding main classes, except for a slightly fainter limit on the observed $r$-band magnitude, $r_{\rm{obs}}<20.1$
  • WISE_VAR_QSO: QSO selection based on variability in WISE light curves. Used the combine band W for variable selection:
    • $N_{epoch}>=9$
    • $A>0.001$
    • $(\gamma>-28.0*A+1.5)$
    • $\chi^{2}>1.0$
    where A is the amplitude of the light curves, $\gamma$ is the time duration exponent, and $\chi^{2}$ is the $\chi^{2}$ of the lightcurves compared to a flat distribution.
  • Z5_QSO: Proposed targets for DESI observations to carry out a systematic survey of $z\sim 5-6.5$ quasars to probe both the evolution of early super-massive black holes (SMBHs) and the cosmic reionization history. The selection using photometric data from only the legacy imaging survey has already discovered 7 quasars at $4.4<z<5.2$ during SV. By adding i band photometry from Pan-STARR1 (PS1), the updated selection will improve the successful rate without significant lost of completeness. PS1 y data and J band data from public NIR surveys are also used to further reject contaminants. The updated selection is based on the DELS DR9, PS1, DR1 & DR2, and public NIR J (e.g., UHS, ULAS, and VHS) data. We build up selection criteria for quasars at redshift ranges of $z\sim 4.8-5.3,5.7-6.4$, and 6.46.86.4-6.8, respectively. The main color-color selections are the (ri)/(iz)(r-i)/(i-z) for $z\sim 5$ and 6, the (iz)/(zy)(i-z)/(z-y) for $z\sim 6$ and 6.5, and z/yW1/W1W2z/y-W1/W1-W2 for all redshifts. Within the DESI footprint, we obtain $\sim 3,500$ targets for $z\sim 5$ quasar, 3,000 targets from $z\sim 6$ selection, and 450 targets for $z>6.4$ quasars, down to a depth of z band magnitude 21.4 for $z\sim 5$ quasars and 21.5 for $z\sim 6$ and 6.5 quasars. We test our selections with both known quasar sample and simulation and estimate the completeness as $60-90\%$ for different redshifts.
  • MWS_RR_LYRAE: We target Gaia DR2 sources with magnitudes $14<G<19$ that are classified as RR Lyrae by the general variability classifier and the Special Object Studies classifier.
  • UNWISE_BLUE/GREEN:
    unWISE catalogs are matched against HSC catalogs with a 2.75" match radius (1 WISE pixel). The unWISE samples are defined in 1901.03337 and 1909.07412. They consist of sources passing a color cut in the WISE bands, and not matched to Gaia stars. Additionally we remove ghosts, latents, and diffraction spikes, and mask around bright infrared stars. For the secondary call, we consider only the green and blue samples. We require deep optical photometry to contain all of the unWISE sources.
    By matching unWISE sources to deep imaging in COSMOS, we find that nearly all of the green (blue) sample has $y<24$ (22.5). Therefore, the y-band depth of the HSC imaging must reach 23.7 for point sources (accounting for the fact that extended source depths are typically 0.3 magnitudes fainter). We make a healpix map of the y magnitude PSF depth from HSC’s “patch_qa" files. However, we find that there is still significant incompleteness towards the edge of each HSC region, where the y-band limiting magnitude drops. Therefore, we also manually define an allowed region for each of the 7 HSC regions that contains HSC imaging to sufficient depth.
    We also use the HSC randoms to remove all regions near a bright star in HSC and require that the randoms are "primary" sources. We also provide an optional flag indicating that a source lies within an nside=8192 healpix pixel with less than 50% of the median number of randoms (i.e. suggesting that it is near a moderately bright star that decreases the number density of randoms). We do not remove sources lying in these regions, as this would reduce the catalog size by $\sim 20\%$, but do flag them for caution in later analyses.Within the green sample, we find 98.3-98.5% of unWISE sources are matched to an HSC source after these cuts. Completeness in blue is 99.3-99.5%. The non-matches are distributed randomly throughout the footprint, and visual inspection reveals that they all either lack a nearby HSC source of sufficient brightness or are weird masking edge cases (i.e. where the unWISE source is unmasked but the nearby HSC source falls in a different healpix pixel and thus is masked).
    Finally, because of the large match radius required by the broad WISE pixels, WISE sources often have multiple HSC matches. To help reduce the number of spurious matches, we remove all HSC sources with $y>24$ in green (22.5 in blue) (this cut is determined by COSMOS matches of unWISE sources, which would suffer from the same spurious-match problem as the HSC sources, except that we can additionally use deep Spitzer imaging to exclude sources where the Spitzer flux is too faint, i.e. they are unlikely to contribute much to the unWISE flux). After this cut, for unWISE sources with multiple matches, we randomly select with probability $exp(-0.5*(\delta/2.72)^{2})$, where Delta is the positional offset between unWISE and HSC, and 2.72" is the 1-sigma width of the WISE PSF. That is, we randomly select the HSC objects with a probability corresponding to the probability of the position scattering by a specified amount just due to the unWISE PSF.
  • WD_BINARIES_BRIGHT:
    These targets are for the White Dwarf binary survey have been selected as follows by cross-matching Gaia and GALEX GR6 and applying the following three cuts: (1) require a 5-sigma parallax detection: parallax/parallax_error >> 5 and (2) select targets in a FUV vs FUV-G color magnitude diagram that lie below the main sequence: $FUV_{mag}+5\cdot log10(parallax/1000)+5>-0.3+1.5\cdot(FUV_{mag}-phot\_g\_mean\_mag)$ and (3) visible for DESI: dec >> -20
    This results in 164,503 targets. Of these, $\sim$ 34,000 overlap with the single white dwarfs that DESI will target for flux calibration purposes, and as science targets in the bright time MWS observations, so that there are $\sim$130,000 net targets for this project.
  • PV_BRIGHT/PV_DARK:
    DESI aims to produce the largest catalog of peculiar velocities ever assembled. This sample will augment constraints on the growth rate of structure from the DESI BGS, reducing the uncertainty by a factor of $\sim 2.5$ at $z<0.1$ compared to the BGS selection and numbers is based on a combination of DR8 and the Siena Galaxy Atlas (SGA), an angular diameter-limited catalog of objects within the DESI footprint that have diameters $D(25)>20\textquotedblright$ (the diameter of the surface brightness isophote at 25 mag/arcsec2 in the $r$-band. Bright and dark delineate if the object was observed during bright or dark time.
  • GC_BRIGHT/DARK
    A target selection of Milky Way globular clusters (GCs). Here, we selected all the high probable member stars (P>0.3>0.3) (Vasiliev & Baumgardt, 2021). Since the selection in this work is based on Gaia EDR3, some GC members might be outside the DESI footprint and may not in LS DR9.
    Then we break the target list into bright time and dark time survey
    • Bright: $16<r_{0}<20$ for bright time survey
    • Dark: $19<r_{0}<21$ for dark time survey
  • TOO_HIP
    These targets are attributed to spare fibers when spare fibers are available.
Special Program Targets: The following targets are part of various spare fiber programs with a brief description of their target selection outlined below. These targets are flagged as TOO_HIP bit in the scnd mask and can be found in the TERTIARY_TARGET column.
  • TERTIARY 1: High Density Low-Z
    The PRINCIPAL sample is a simple magnitude-limited sample of $z_{\rm{fiber}}<21.6$ with a stellar rejection cut in rzr-z and zw1z-w1 (similar to the DESI LRGs, although slightly relaxed to accommodate lower redshift galaxies):
    • $(z-w1)>0.8\cdot(r-z)-1.1$
    Unlike the DESI LRG selection, we also remove objects with extended morphology (i.e., objects that are not “PSF” type) from the PRINCIPAL sample.
    The filler (FILLER_HIP,FILLER_LOP) sample extends the magnitude limit to $z_{\rm{fiber}}<22.4$, and we use color cuts in grg-r and rzr-z to assign a higher priority to ($z>\sim 0.7$) galaxies (which are roughly half of the $21.6\leq z_{\rm{fiber}}<22.4$ target sample; few lower-priority filler targets were assigned fibers):
    • $(g-r)<1.2\cdot(r-z)$
    • (rz)>1.3(r-z)>1.3
    For both the principal and filler samples, we exclude from them objects that have already been observed by DESI with DELTACHI2 >40>40. We also excluded SV3 BGS targets with $r_{\rm{fiber}}<21.0$ (fainter BGS galaxies are retained here because they have lower redshift success rates in SV3) and any SV3 LRG and QSO targets as we already observed them with high completeness and high success rates in SV3.
    The combined sample is observed with 11 DESI tiles in a “rosette” pattern in the COSMOS field centered at $R.A.=150.100^{\circ},Dec.=2.182^{\circ}$. One 1200s (effective time) exposure is requested for each PRINCIPAL target. Four 1200s exposures are requested for each filler target (although only a small fraction of them were observed to full depth with the 11 tiles).
  • TERTIARY 5/6/7/8: Calibration These fields are ordinary dark and bright tiles that get repeatedly re-observed over the course of the survey. We observe one dark and bright calibration field every lunation. On each field location, we design a series of tiles with slight offsets from one another so that the same targets can be imaged on different fibers.
  • TERTIARY 15: SN-HOSTS/DESI_DEEP_WL/ELG
    DESI_DEEP: We selected objects with extinction-corrected $i$-band magnitudes between 22 and 24.5 and $i$-fiber magnitude between 22 and 25 for the COSMOS field and 22 and 24.75 for the XMM-LSS field from the HSC-PDR3 wide catalog located. The faint limit was set to be similar to the depth of LSST Year-1 data, and the HSC weak lensing sample. The number of objects was then subsampled to have a uniform distribution of $i$-band magnitudes. The targets were then divided into two bins of priority, more objects with $i$-band magnitude between 22 and 23 were observed but for shorter exposure times whereas a smaller number of objects with $i$-band magnitude between 23 and 24.5 were observed but with larger exposure time. Further details on the target selection can be found in Dey et al. 2025 (in prep).
    SN_HOSTS: Our candidates are selected by the presence of a SN in the host galaxy within the HSC COSMOS deep field (Yasuda et al., 2019). There is no magnitude or color cut on the targeted host but we cut at $r=25$ for the host galaxies of less certain SNe. Internally the targets are prioritized according to their likely scientific value for constructing a Hubble diagram and measuring the dark energy equation of state. Highest priority went to host galaxies of SNeIa observed with HST as part of programs GO14808 and GO15363 (PI Suzuki); priority then decreased in the order of: other likely SNe Ia; including the HST SNe Ia), possible SNe Ia, and less certain SNe Ia. We then removed the targets whose redshifts we had previously obtained with AAOmega Sharp et al. (2006), and targets with redshifts from the literature. To make a more complete Hubble diagram, we also include host galaxies for confirmed or suspected SNe Ia lacking redshifts with the SNLS D2 deep field Astier et al. (2006).
  • TERTIARY 19: DESI II Bright
    We propose to observe $\sim$4000 galaxies in the XMM-LSS field in $i\in[19,22]$ during typical or favorable bright time conditions, fairly drawn from a parent target list with density 3125 per square degree to fill the available number of fibers (resulting in a final density of $\sim$480 per square degree). We propose 4000 sec of effective time, with dithering. Our selection observes objects at $i>22$. We estimate the necessary observing time based on the “filler hip” high density sample from the COSMOS field, which peaks at $i\sim 22$ and achieved $99\%$ redshift success rate with 4800s of effective time. Similarly to that sample, we will apply a cut in $i$-band fiber magnitude to remove objects with very low surface brightness that are very unlikely to have successful redshifts after our full integration time. We also subsample galaxies with an analytic exponential weight such that they are roughly uniformly distributed in magnitude.
  • TERTIARY 23: 4-in-1
    LOWZ: See LOWZ
    MERIAN: The Merian Ancillary Program consists of 3 selections
    • MS1: dwarf galaxies with $7.8<logM_{star}/M_{sun}<9.5$ at $z<0.25$ and $19.5<i<23$. The mass and redshift for this sample is based on the COSMOS2015 catalog 30-band SED fitting.
    • MS2: all galaxies with $i<25$ at $z<0.12$ based on Merian 7-band photo-z’s.
    DESI_DEEP: See TERTIARY 15
    ELG: Two extended selection of ELGs are a simple grizy cut based on HSC/WIDE
    1. 1.
      Selection 1:
      • $g_{cmodel\_mag}-a_{g}>22.0$
      • $g_{cmodel\_mag}-a_{g}<24.5$
      • $r_{cmodel\_mag}-a_{r}-i_{cmodel\_mag}+a_{i}>0.1$
      • $r_{cmodel\_mag}-a_{r}-i_{cmodel\_mag}+a_{i}<0.4$
      • $g_{cmodel\_mag}-a_{g}-r_{cmodel\_mag}+a_{r}>-0.1+0.7\cdot(r_{cmodel\_mag}-a_{r}-i_{cmodel\_mag}+a_{i})$
      • $g_{cmodel\_mag}-a_{g}-r_{cmodel\_mag}+a_{r}<0.1+0.7\cdot(r_{cmodel\_mag}-a_{r}-i_{cmodel\_mag}+a_{i})$
      • $z_{cmodel\_mag}-a_{z}-y_{cmodel\_mag}+a_{y}>0.35$
    2. 2.
      Selection 2:
      • (iy>0.35i-y>0.35) & (ri<iy0.19)(r-i<i-y-0.19)
      • ($g<24$ & $g_{\rm{fiber}}<24.3$) OR $(r<24$ & $r_{\rm{fiber}}<24.3)$
  • TERTIARY 26: ELG (Special)
    Selection of Special Program ELGs(see TERTIARY 23 ELGs) that received less than 4 observations.
  • TERTIARY 38: ELG Systematics
    This sample is an expanded ELG selection for the use of forward-modeling studies of the ELG imaging systematics (Zhou et al., 2025). The imaging data comes from DECam deep field imaging and HSC PDR3 deep/ultra-deep imaging in COSMOS.
    DECam selection:
    • $19.5<\texttt{gfibermag}<24.5$
    • gr<0.9g-r<0.9
    • $g-r<1.32-0.7\cdot(r-z)$
    HSC selection:
    • $19.5<\texttt{gfibermag}<24.55$
    • gr<0.8g-r<0.8
    • $g-r<1.22-0.7\cdot(r-z)$
    where gfibermag is based on g_apertureflux_15_flux. The combined target sample (excluding duplicates from the two selections) has a density of roughly 24000/sq. deg. It is observed by 25 tiles in a “rosette" pattern in the COSMOS field centered at $R.A.=150.11^{\circ},Dec.=2.3^{\circ}$, each tile with 1000s effective exposure time. A total of 97800 unique objects in the target sample were observed.
Appendix B Targeting Information
The tables below summarize all the target classes within each catalog, the number of quality redshifts, and the fraction of quality redshifts for that class. Note that objects in these tables are represented by their target selection, not by spectroscopic classification. As there is overlap between galaxy, star and quasar spectroscopic classification within a target class, numbers for total objects and quality redshift may appear higher than previously mentioned. Targets may share multiple target classes as their photometric properties satisfy multiple selection algorithms.
Table 8: COSMOS Target Classification and Redshift Robustness
Target Class Total Objects Quality Redshifts Quality Redshift (%)
LRG 11702 11510 98.4%
ELG 35522 29468 83.0%
QSO 4868 4374 89.9%
BGS_ANY 21691 21326 98.3%
MWS_ANY 17905 17866 99.8%
STD_FAINT 1106 1101 99.5%
STD_BRIGHT 564 561 99.5%
STD_WD 92 89 96.7%
SCND_ANY 156243 122237 78.2%
Secondary Targets
DARK_TOO_HIP 150802 117677 78.0%
LOW_Z_TIER3 1915 1864 97.3%
WISE_VAR_QSO 1593 1482 93.0%
PSF_OUT_DARK 1405 1307 93.0%
LOW_Z 886 851 96.0%
PV_DARK_HIGH 813 520 64.0%
PV_BRIGHT_HIGH 736 443 60.2%
HSC_HIZ_SNE 753 625 83.0%
BRIGHT_TOO_HIP 723 596 82.4%
LOW_Z_TIER2 706 665 94.2%
PSF_OUT_BRIGHT 645 622 96.4%
MWS_FAINT_BLUE 612 606 99.0%
UNWISE_BLUE 306 301 98.4%
PV_DARK_MEDIUM 208 199 95.7%
PV_BRIGHT_MEDIUM 201 198 98.5%
LOW_Z_TIER1 168 146 86.9%
PV_DARK_LOW 123 117 95.1%
PV_BRIGHT_LOW 97 90 92.8%
MWS_FAINT_RED 64 64 100.0%
ISM_CGM_QGP 60 55 91.7%
UNWISE_GREEN 56 55 98.2%
QSO_RED 27 23 85.2%
MWS_CALIB 26 26 100.0%
BACKUP_CALIB 26 26 100.0%
FAINT_HPM 25 20 80.0%
SN_HOSTS 21 21 100.0%
STRONG_LENS 22 17 77.3%
MWS_CLUS_GAL_DEEP 17 14 82.4%
MWS_RR_LYRAE 6 6 100.0%
WD_BINARIES_BRIGHT 6 6 100.0%
HPM_SOUM 5 4 80.0%
BHB 3 3 100.0%
LOW_MASS_AGN 3 3 100.0%
Z5_QSO 2 2 100.0%
MWS_RRLYR 1 1 100.0%
Special Targets
ELG_PRINCIPAL 81583 60766 74.5%
PRINCIPAL 33083 32278 97.6%
ELG 19007 13908 73.2%
FILLER_HIP 5270 4068 77.2%
LOWZ_FAINT 2808 1823 64.9%
BGS_BGS_BRIGHT 1809 1784 98.6%
FILLER_LOP 1757 1684 95.8%
DESI_LRG 1115 1109 99.5%
MERIAN_MS2 1026 539 52.5%
MERIAN_MS1 1004 858 85.5%
DESI_DEEP_HIP 932 656 70.4%
BGS_BGS_FAINT 689 683 99.1%
DESI_ELG_LOP 666 554 83.2%
DESI_QSO 627 572 91.2%
LOWZ_BRIGHT 593 569 96.0%
DESI_DEEP_LOP 577 484 83.9%
DESI_ELG_HIP 479 422 88.1%
BGS_BGS_FAINT_HIP 267 266 99.6%
MWS_MWS_MAIN_BLUE 266 265 99.6%
MWS_MWS_BROAD 192 191 99.5%
DESI_DEEP_HIP2 136 132 97.1%
DESI_ELG 86 86 100.0%
MWS_MWS_MAIN_RED 65 64 98.5%
MWS_MWS_WD 17 16 94.1%
MWS_MWS_BHB 7 7 100.0%
BGS_BGS_WISE 4 4 100.0%
MWS_MWS_NEARBY 1 1 100.0%
Table 9: XMM Target Classifications and Redshift Robustness
Target Class Total Objects Quality Redshifts Quality Redshift (%)
LRG 9986 9738 97.5%
ELG 28058 22429 79.9%
QSO 5253 4714 89.7%
BGS_ANY 18359 18023 98.2%
MWS_ANY 10175 10141 99.7%
SCND_ANY 11222 9525 84.9%
STD_FAINT 820 818 99.8%
STD_BRIGHT 465 463 99.6%
STD_WD 70 69 98.6%
QSO_Z5 25 25 100.0%
Secondary Targets
DARK_TOO_HIP 3364 2422 72.0%
PSF_OUT_DARK 2469 2336 94.6%
BRIGHT_TOO_HIP 1591 1425 89.6%
LOW_Z_TIER3 1488 1425 95.8%
WISE_VAR_QSO 1858 1744 93.9%
PSF_OUT_BRIGHT 720 684 95.0%
PV_DARK_HIGH 353 267 75.6%
LOW_Z_TIER2 258 202 78.3%
PV_BRIGHT_HIGH 319 261 81.8%
LOW_Z_TIER1 140 97 69.3%
PV_DARK_MEDIUM 145 140 96.6%
PV_BRIGHT_MEDIUM 140 140 100.0%
STRONG_LENS 81 64 79.0%
FAINT_HPM 76 19 25.0%
PV_DARK_LOW 55 51 92.7%
QSO_RED 37 29 78.4%
PV_BRIGHT_LOW 51 49 96.1%
MWS_FAINT_BLUE 66 64 97.0%
MWS_FAINT_RED 35 35 100.0%
SN_HOSTS 16 16 100.0%
LOW_MASS_AGN 12 12 100.0%
BHB 4 4 100.0%
WD_BINARIES_BRIGHT 2 2 100.0%
Z5_QSO 2 2 100.0%
MWS_RR_LYRAE 2 2 100.0%
WD_BINARIES_DARK 1 1 100.0%
BRIGHT_HPM 1 1 100.0%
Special Targets
BGS_BGS_BRIGHT 2729 2719 99.6%
DESI_DEEP_LOP 1981 1605 81.0%
DESI_LRG 1891 1877 99.3%
DESI_ELG_LOP 1734 1445 83.3%
DESI_QSO 1584 1436 90.7%
PRINCIPAL 1179 1140 96.7%
FILLER 1152 1142 99.1%
DESI_DEEP_HIP 1025 624 60.9%
DESI_ELG_HIP 839 748 89.2%
BGS_BGS_FAINT 701 699 99.7%
MWS_MWS_BROAD 434 434 100.0%
BGS_BGS_FAINT_HIP 421 418 99.3%
MWS_MWS_MAIN_BLUE 362 362 100.0%
FILLER_LOP 340 324 95.3%
SNHOST 286 188 65.7%
DESI_ELG 149 145 97.3%
SNHOST_HIP 124 92 74.2%
MWS_MWS_MAIN_RED 104 104 100.0%
MWS_MWS_WD 18 17 94.4%
SNHOST_VERY_HIP 10 9 90.0%
BGS_BGS_WISE 9 9 100.0%
MWS_MWS_BHB 4 4 100.0%
MWS_MWS_NEARBY 1 1 100.0%
[…]
Appendix C Data Model
Column Name Unit Dtype Description
DESINAME |S22 Unique DESI target identifier
TARGET_RA deg >f8>f8 Barycentric right ascension in ICRS
TARGET_DEC deg >f8>f8 Barycentric declination in ICRS
OBJTYPE |S3 Object type: TGT, SKY, NON, BAD
PMRA mas / yr >f4>f4 Proper motion in the +RA direction (already including cos(dec))
PMDEC mas / yr >f4>f4 Proper motion in the +Dec direction
SURVEY |S7 Survey when the object’s best spectrum was observed (Main, SV1, SV3, Special)
PROGRAM |S6 Program when the object’s best spectrum was observed (bright, dark, backup, other)
HEALPIX >i4>i4 HealPIX pixel of the object’s best spectrum
Z >f8>f8 Redshift estimation from Redrock or Redrock+afterburners for quasar targets for the object’s best spectrum
ZERR >f8>f8 Redshift uncertainty estimation from Redrock or Redrock+afterburners for quasar targets for the object’s best spectrum
ZWARN >i8>i8 Warning flag bit for Redrock for the object’s best spectrum; 0 indicates no issue; 4 indicates low DELTACHI2
DELTACHI2 >f8>f8 Difference in $\chi^{2}$ from Redrock’s best PCA template fit and second-best fit for the object’s best spectrum (used for quality assessments)
CHI2 >f8>f8 Non-reduced $\chi^{2}$ from Redrock’s best PCA template fit for the object’s best spectrum
NPIXELS >i8>i8 Number of unmasked pixels contributing to the Redrock fit
SPECTYPE |S6 Spectral classification from Redrock or Redrock+afterburners for quasar targets for the object’s best spectrum
SUBTYPE |S3 Spectral subtype of the object for the object’s best spectrum (may be blank)
BEST_Z >f8>f8 Best reported redshift for this object
DZ >f8>f8 Largest difference in redshifts between objects that were merged
QUALITY_Z bool Flag (0 or 1) indicating if the object has a quality redshift
TSNR2_BGS >f4>f4 Target signal-to-noise ratio squared for Bright Galaxy Survey targets for the best spectrum
TSNR2_LRG >f4>f4 Target signal-to-noise ratio squared for Luminous Red Galaxy targets for the best spectrum
TSNR2_ELG >f4>f4 Target signal-to-noise ratio squared for Emission Line Galaxy targets for the best spectrum
TSNR2_QSO >f4>f4 Target signal-to-noise ratio squared for Quasar targets for the best spectrum
TSNR2_LYA >f4>f4 Target signal-to-noise ratio squared for Lyman-alpha targets for the best spectrum
IS_QSO_MGII bool Indicates if a quasar target was flagged as a quasar by the MgII afterburner for the object’s best spectrum
A >f4>f4 Fitted parameter by MgII
B >f4>f4 Fitted parameter by MgII
Z_NEW >f8>f8 New redshift computed with Redrock with QN prior and only QSO templates
ZERR_NEW >f4>f4 Redshift error from the new run of Redrock
Z_QN >f4>f4 Redshift computed with Quasarnp
IS_QSO_QN_NEW_RR bool Indicates if the object is detected as a QSO with Quasarnp and a new redshift fit with prior is performed for the object’s best spectrum
DESI_TARGET >i8>i8 DESI (dark time program) target selection bitmask
BGS_TARGET >i8>i8 BGS (Bright Galaxy Survey) target selection bitmask
MWS_TARGET >i8>i8 Milky Way Survey targeting bits
SCND_TARGET >i8>i8 Target selection bitmask for secondary programs
SV1_DESI_TARGET >i8>i8 DESI (dark time program) target selection bitmask for SV1
SV1_BGS_TARGET >i8>i8 BGS (Bright Galaxy Survey) target selection bitmask for SV1
SV1_MWS_TARGET >i8>i8 Milky Way Survey targeting bits for SV1
SV1_SCND_TARGET >i8>i8 Target selection bitmask for secondary programs for SV1
SV3_DESI_TARGET >i8>i8 DESI (dark time program) target selection bitmask for SV3
SV3_BGS_TARGET >i8>i8 BGS (Bright Galaxy Survey) target selection bitmask for SV3
SV3_MWS_TARGET >i8>i8 Milky Way Survey targeting bits for SV3
SV3_SCND_TARGET >i8>i8 Target selection bitmask for secondary programs for SV3
SPECIAL_TARGET |S100 Target selection description for special programs
LINENAME_FLUX $1\times 10^{-17}\mathrm{erg}/(\mathrm{s}\,\mathrm{cm}^{2})$ >f4>f4 Gaussian-integrated emission-line flux measurement from FastSpecFit of the best spectrum
LINENAME_FLUX_IVAR $1\times 10^{34}\mathrm{cm}^{4}\mathrm{s}^{2}/\mathrm{erg}^{2}$ >f4>f4 Inverse variance of integrated flux measurement from FastSpecFit of the best spectrum
TOTAL_NUM_COADD >f8>f8 Number of exposures used in the coadded spectrum for the object’s best spectrum
TOTAL_COADD_EXPTIME s >f8>f8 Total exposure time across all coadded exposures
N_EPOCHS Number of different observational epochs for the target
LRG_MASK >f8>f8 Bright star mask used for LRG targets applied to the catalog
ELG_MASK >f8>f8 Bright star mask used for ELG targets applied to the catalog
HAS_RVS bool Boolean flag indicating RVS measurements for object
VRAD km / s >f8>f8 Radial velocity
VRAD_ERR km / s >f8>f8 Radial velocity error
VRAD_SKEW >f8>f8 Radial velocity skewness
VRAD_KURT >f8>f8 Radial velocity kurtosis
LOGG >f8>f8 Log of surface gravity
TEFF K >f8>f8 Effective temperature
FE_H >f8>f8 [Fe/H] from template fitting
LOGG_ERR >f8>f8 Log of surface gravity uncertainty
TEFF_ERR K >f8>f8 Effective temperature uncertainty
FE_H_ERR >f8>f8 [Fe/H] from template fitting uncertainty
VSINI km / s >f8>f8 Stellar rotation velocity
RVS_WARN >i8>i8 RVSpecfit warning flag
HAS_DECALS_DR9 bool Flag indicating if the object has corresponding photometry from DECaLS
HAS_DECAM_DR10 bool Flag indicating if the object has corresponding photometry from DECam
HAS_HSC_UD_PDR3 bool Flag indicating if the object has corresponding photometry from HSC Ultra Deep
HAS_HSC_WIDE_PDR3 bool Flag indicating if the object has corresponding photometry from HSC WIDE
HAS_COSMOS2020 bool Flag indicating if the object has corresponding photometry from COSMOS2020
HAS_MERIAN bool Flag indicating if the object has corresponding photometry from Merian
HAS_VI_Z bool Boolean flag indicating if object has visually inspected redshift
VI_Z >f8>f8 Visually inspected redshift
VI_SPECTYPE |S6 Visually inspected spectroscopic classification
VI_QUALITY >f8>f8 Visually inspected redshift quality (3 and 4 indicate quality visually inspected redshift)